### v1

# Optimal Control of an SIR Model with Noncompliance as a Social Contagion

Abstract.
We propose and study a compartmental model for epidemiology with human behavioral effects. Specifically, our model incorporates governmental prevention measures aimed at lowering the disease infection rate, but we split the population into those who comply with the measures and those who do not comply and therefore do not receive the reduction in infectivity. We then allow the attitude of noncompliance to spread as a social contagion parallel to the disease. We derive the reproductive ratio for our model and provide stability analysis for the disease-free equilibria. We then propose a control scenario wherein a policy-maker with access to control variables representing disease prevention mandates, treatment efforts, and educational campaigns aimed at encouraging compliance minimizes a cost functional incorporating several cost concerns. We characterize optimal controls via the Pontryagin optimality principle and present simulations which demonstrate the behavior of the control maps in several different parameter regimes.
1. Introduction
In this paper, we are interested in incorporating human behavioral effects into a Susceptible-Infected-Recovered (SIR) type compartmental model for epidemic spread of a generic infectious disease. At the outset of an epidemic, governments will enact public health protocols meant to slow disease spread. However, adherence with non-pharmaceutical intervention (NPI) measures such as mask-wearing, shelter-at-home mandates, and social distancing is certainly imperfect. For example, Dolan provides an interesting account of opposition to prolonged public health protocols during the 1918 Flu epidemic and the actions of the so-called “anti-mask league” [20], and in a study by the Pew Research Center from June 4-10, 2020 during the height of the COVID 19 pandemic, 16% of $9654$ Americans surveyed reported that they “hardly ever” or “never” wear a mask [23]. Likewise, the Angus Reid Institute reported on July 15, 2020 that 45% of 1503 Canadians surveyed responded that they “rarely” or “never” wear masks, and among the 1202 participants who reported that they do not always wear masks, 15% reported that their main reason for not wearing a mask is that they believe masks are ineffective and 11% reported their main reason for not wearing a mask as “no one else is wearing them” [1]. Other studies reported similar levels of nonadherence with several types of NPIs [5, 10, 33]. From reports like this, it appears that there will always be nontrivial subpopulations who do not comply with NPI measures, and given evidence that NPI measures have a positive effect on epidemic outcomes [30], accounting for this noncompliance is key to increasing fidelity in mathematical models of epidemiology.
Accordingly, we propose a compartmental model for epidemiology wherein the population is split into those who comply with public health protocols (dubbed the compliant population), and those who do not comply (dubbed the noncompliant population). The disease is assumed to spread slower among the compliant population. Social contagion theory [12] posits that attitudes, opinions and behaviors can spread among a population in a manner similar to that of an infectious disease, and has been used to model prescription and illicit drug use [11, 16, 22], alcohol consumption [48], governmental corruption [32], adverse mental health conditions such as depression [40], and public sentiment [31] in a variety of contexts including opinions regarding anti-smoking legislation [35] and, importantly, vaccine and treatment uptake during epidemics [9, 24]. Given this, and the aforementioned study wherein people reported not wearing masks because “no one else is wearing them” [1], we suppose that the attitude of noncompliance with NPIs spreads throughout the population as a social contagion. We further imagine a policy-maker with several different options for epidemic control who aims to minimize a cost functional which trades off costs for increased infections and increased noncompliance with economic costs for increased use of controls.
Incorporation and assessment of the effects of heterogeneous behavior into epidemiological models is becoming increasingly popular (see, for example, [4, 21, 27, 34, 36], or [43] where the authors propose a rather general framework similar to the modeling we use below, though with linear transfer between behavior classes). Likewise, there are several recent studies of optimal control for epidemiological models in different modeling scenarios, such as optimal treatment or vaccination [6, 13, 26, 50], optimal quarantine under a variety of limitations [3, 15], and assessment of the effects of contact heterogeneity [14] or the use of different types of cost functionals [44]. However, very few studies include both optimal control for epidemiology and mechanistic modeling of human behavior. One exception is [8] which takes a more theoretical look at a system of controlled partial differential equations describing an epidemic. Other studies incorporate control and human behavior more implicitly, by, for example, allowing infection rates to depend on ambient dynamics under the assumption that people will change their behavior based on local disease effects [37] or media coverage [18].
In this manuscript, we develop a compartmental epidemiological model which incorporates human behavior in an explicit, mechanistic manner. We then consider optimal control of the model and observe the differences in optimal control strategies in different parameter regimes. The remainder of the manuscript is organized as follows. In section 2, we introduce our model, derive the reproductive ratio, and discuss stability of disease-free equilibria (though a full proof of our stability theorem is relegated to appendix A). In section 3, we introduce the optimal control problem that we will solve and give a characterization of the optimal control maps via the Pontryagin optimality principle. In section 4, we discuss the results of simulation of our model in different parameter regimes. Finally, section 5 includes a brief conclusion and discussion of avenues for future work.
2. An SIR Model with Behavioral Effects: Disease-Free Equilibria and Reproductive Ratios
As with all ordinary and partial differential equation (O/PDE) based epidemiological models, we build off of the classical SIR model of Kermack and McKendrick [25]. We assume a constant birth rate $b>0$ into the susceptible population, and that deaths occur proportional to population sizes with natural rate $\delta>0$. The infection spreads via mass action with rate $\beta>0$ and infectious individuals recover with rate $\gamma>0$. This leads to the SIR model
(2.1) $\begin{split}\frac{dS}{dt}&=b-\beta SI-\delta S,\\ \frac{dI}{dt}&=\beta SI-(\gamma+\delta)I,\\ \frac{dR}{dt}&=\gamma I-\delta R.\end{split}$
As mentioned above, we incorporate human behavior in a manner inspired by social contagion theory. Each population is divided into a compliant portion (denoted by $S,I,R$) and a noncompliant portion (denoted by $S^{*},I^{*},R^{*}$), and we assume that, parallel to the disease, noncompliance spreads via a second mass action term. With this assumption, we propose the following model:
(2.2) $\begin{split}\frac{dS}{dt}&=(1-\xi)b-\beta(1-\alpha(t))S(I+I^{*})-(\overline{\mu}-\mu(t))SN^{*}+\nu(t)S^{*}-\delta S,\\ \frac{dI}{dt}&=\beta(1-\alpha(t))S(I+I^{*})-(\gamma+\eta(t))I-(\overline{\mu}-\mu(t))IN^{*}+\nu(t)I^{*}-\delta I,\\ \frac{dR}{dt}&=(\gamma+\eta(t))I-(\overline{\mu}-\mu(t))RN^{*}+\nu(t)R^{*}-\delta R,\\ \frac{dS^{*}}{dt}&=\xi b-\beta S^{*}(I+I^{*})+(\overline{\mu}-\mu(t))SN^{*}-\nu(t)S^{*}-\delta S^{*},\\ \frac{dI^{*}}{dt}&=\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu(t))IN^{*}-\nu(t)I^{*}-\delta I^{*},\\ \frac{dR^{*}}{dt}&=\gamma I^{*}+(\overline{\mu}-\mu(t))RN^{*}-\nu(t)R^{*}-\delta R^{*}.\end{split}$
Here the birth rate $b$, death rate $\delta$, infectivity rate $\beta$ and recovery rate $\gamma$ have the same interpretation as in (2.1). New parameters are as follows: $\xi\in[0,1]$ is the probability with which newly introduced members of the population are noncompliant, and hence the birth rate $b$ is split between the $S$ and $S^{*}$ equations with contributions $(1-\xi)b$ and $\xi b$ respectively. The total noncompliant population is defined by $N^{*}=S^{*}+I^{*}+R^{*}$ and is involved in new mass action terms which cause transfer from $S,I,R$ to $S^{*},I^{*},R^{*}$ with baseline infectivity rate $\overline{\mu}$ (the assumption being that anytime a compliant individual comes in contact with a noncompliant individual, they have some chance of becoming noncompliant). The remaining parameters $\alpha,\eta,\mu,\nu$ represent the policy-maker’s attempts to stop the spread of the disease and/or noncompliance. While we will typically suppress their $t$-dependence, we have written them as functions of $t$ in (2.2) to emphasize that these will be treated as control variables that the policy-maker can adjust throughout the course of the epidemic. Their respective interpretations and admissible values are:
  • (1)
    $\alpha(\cdot)\in[0,1]$ - reduction in infectivity among interactions with the compliant susceptible population.
  • (2)
    $\eta(\cdot)\in[0,\overline{\eta}]$ - increased recovery rate among compliant infectious population due to treatment seeking. Here $\overline{\eta}$ is the maximum achievable increase in recovery rate for the disease due to treatment.
  • (3)
    $\mu(\cdot)\in[0,\overline{\mu}]$ - reduction in spread of noncompliance due to public health information initiatives.
  • (4)
    $\nu(\cdot)\in[0,\overline{\nu}]$ - rate of recovery from noncompliance due to educational campaigns aimed at increasing compliance with NPIs. Here $\overline{\nu}$ is the maximum achievable recovery rate from noncompliance.
The uncontrolled scenario would be $\alpha(\cdot),\eta(\cdot),\mu(\cdot),\nu(\cdot)\equiv 0$, and the maximally controlled scenario would have each control variable identically equal to its maximum value. Note that we do not consider vaccination, which can be modeled as either an ordinary or “singular” control variable as discussed in [26], opting instead to focus on the early stages of infection when vaccination is unavailable and NPIs play a larger role. In this way, our model is similar to the ones analyzed in [7] in the ODE setting, [39] in the stochastic ODE setting, and [8, 38] in the PDE setting. However, [7, 38, 39] do not consider the optimal control problem, and the work in [8] does not incorporate treatment as a control, and is also more focused on analytic results regarding existence and characterization of solutions to a PDE-constrained optimization problem. Here we are interested in empirically observing features of the optimal control maps in different parameter regimes.
In the ensuing section, we will introduce a cost functional for the policy-maker, and evaluate optimal control strategies. Before that, we cover some standard considerations regarding compartmental epidemiological models; namely, nonnegativity of solutions, derivation of the reproductive ratio, and stability of disease-free equilibria. For the remainder of this section, we treat $\alpha,\eta,\mu,\nu$ as fixed, nonnegative, constant values. We will go through this quickly; much of the same work is included in more detail in [39], albeit for a slightly different system of equations.
Because we will use the next generation matrix method [19, 46, 47] to derive the reproductive ratio for our system, we write (2.1) abstractly as
(2.3) $x^{\prime}=\mathcal{F}(x)+\mathcal{V}^{+}(x)-\mathcal{V}^{-}(x)$
where we reorder the variables $x=(I,I^{*},S,R,S^{*},R^{*})$ by bringing the infected compartments to the front, and define
(2.4) $\mathcal{F}(x)=\begin{bmatrix}\beta(1-\alpha)S(I+I^{*})\\ \beta S^{*}(I+I^{*})\\ 0\\ 0\\ 0\\ 0\end{bmatrix}$
and
(2.5) $\mathcal{V}^{+}(x)=\begin{bmatrix}\nu I^{*}\\ (\overline{\mu}-\mu)IN^{*}\\ (1-\xi)b+\nu S^{*}\\ (\gamma+\eta)I+\nu R^{*}\\ \xi b+(\overline{\mu}-\mu)SN^{*}\\ \gamma I^{*}+(\overline{\mu}-\mu)RN^{*}\end{bmatrix},\,\,\,\,\,\,\,\,\,\,\mathcal{V}^{-}(x)=\begin{bmatrix}(\gamma+\eta+\delta+(\overline{\mu}-\mu)N^{*})I\\ (\gamma+\nu+\delta)I^{*}\\ (\beta(1-\alpha)(I+I^{*})+(\overline{\mu}-\mu)N^{*}+\delta)S\\ ((\overline{\mu}-\mu)N^{*}+\delta)R\\ (\beta(I+I^{*})+\nu+\delta)S^{*}\\ (\nu+\delta)R^{*}\end{bmatrix}$.
Here $\mathcal{F}$ includes all terms which correspond to adding new infections into the system, $\mathcal{V}^{+}$ includes all terms corresponding to in-flow to a compartment caused by anything other than new infections, and $\mathcal{V}^{-}$ includes all terms corresponding to outflow from a compartment for any reason. We note in particular that $\mathcal{F},\mathcal{V}^{+},\mathcal{V}^{-}\geq 0$ when $S,I,R,S^{*},I^{*},R^{*}\geq 0$.
Proposition 2.1.
Assuming nonnegative initial conditions, (2.2) admits a unique, global-in-time, nonnegative solution.
Proof.
A unique local solution to (2.2) exists as consequence of the Picard-Lindelöf theorem since the nonlinearities are Lipschitz. Nonnegativity is preserved as long as the solution exists because, looking at each component $x_{j}$ of (2.3), we have $x_{j}^{\prime}\geq 0$ when $x_{j}=0$ since the $i^{\text{th}}$ component of $\mathcal{V}^{-}$ is proportional to $x_{i}$. Given nonnegativity, we see that solutions cannot blow up because the total population $N_{\text{total}}=S+I+R+S^{*}+I^{*}+R^{*}$ satisfies
(2.6) $N_{\text{total}}^{\prime}=b-\delta N_{\text{total}}\,\,\,\,\,\implies\,\,\,\,\,N_{\text{total}}(t)=\frac{b}{\delta}+\bigg{(}N_{\text{total}}(0)-\frac{b}{\delta}\bigg{)}e^{-\delta t}\leq\max\left\{N_{\text{total}}(0),\frac{b}{\delta}\right\}$
which serves as an upper bound on each of the subpopulations. ∎
We can assume without loss of generality that $N_{\text{total}}(0)=1$, whereupon alternately setting $\tfrac{b}{\delta}\geq 1$ or $\tfrac{b}{\delta}<1$ would correspond to situations where the population is increasing or where the population is strictly decreasing, respectively. For ease of analysis, we will always assume that $\tfrac{b}{\delta}\geq 1=N(0)$, so that $\tfrac{b}{\delta}$ serves as a bound on the total population.
A disease-free equilibrium (DFE) point for (2.2) is an equilibrium point wherein $I=I^{*}=0$. By adding the equations for $R,R^{*}$ together, we see that any DFE will also have $R=R^{*}=0$, whereupon $x=(0,0,s,0,s^{*},0)$ is a DFE if and only if
(2.7) $\begin{split}(1-\xi)b-(\overline{\mu}-\mu)ss^{*}+\nu s^{*}-\delta s&=0,\\ \xi b+(\overline{\mu}-\mu)ss^{*}-\nu s^{*}-\delta s^{*}&=0.\end{split}$
Because we are looking for physically meaningful solutions, we only consider values $0\leq s,s^{*}\leq\frac{b}{\delta}$. With this restriction in mind, we specify the different parameter regimes which lead to existence of different DFEs. If we assume $\xi=0$ so that all newly introduced members of the population are compliant, there are two solutions which can be physically meaningful:
(2.8) $\begin{split}s_{1}&=\tfrac{b}{\delta},\\ s^{*}_{1}&=0,\end{split}\hskip 56.9055pt\text{ and }\hskip 56.9055pt\begin{split}s_{2}&=\tfrac{\nu+\delta}{\overline{\mu}-\mu},\\ s^{*}_{2}&=\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu},\end{split}$
The first is always physically meaningful and represents the “best case scenario” for the policy-maker wherein the whole population is compliant. The second is physically meaningful and distinct only if $\tfrac{b}{\delta}>\frac{\nu+\delta}{\overline{\mu}-\mu}.$ Since $\mu\in[0,\overline{\mu}]$, the denominator of $\frac{\nu+\delta}{\overline{\mu}-\mu}$ is nonnegative, so in the case that $\mu\to\overline{\mu}$, the second DFE is no longer meaningful.
If instead $\xi\in(0,1]$ so that some portion of the newly introduced population is noncompliant, (2.7) has a unique physically meaningful solution
(2.9) $\begin{split}s_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}+\tfrac{\delta+\nu}{\overline{\mu}-\mu}-\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\overline{\mu}-\mu}\right)^{2}+\tfrac{4\xi b}{\overline{\mu}-\mu}}\right),\\ s^{*}_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\overline{\mu}-\mu}+\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\overline{\mu}-\mu}\right)^{2}+\tfrac{4\xi b}{\overline{\mu}-\mu}}\right).\end{split}$
In the limit as $\xi\to 0$, we find $(s_{3},s_{3}^{*})\to(s_{1},s_{1}^{*})$ when $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\overline{\mu}-\mu}$, and $(s_{3},s_{3}^{*})\to(s_{2},s_{2}^{*})$ when $\tfrac{b}{\delta}>\frac{\nu+\delta}{\overline{\mu}-\mu}$. A special case of (2.9) is the “worst case scenario” for the policy-maker wherein $\xi=1$ and $\nu=0$, meaning that all newly introduced members of the population are noncompliant and noncompliance is a permanent state. In this case, the DFE is $s_{3}=0,s^{*}_{3}=\tfrac{b}{\delta}$.
Given this, we can compute the reproductive ratio $\mathscr{R}_{0}(s,s^{*})$ corresponding to the DFE $x_{s,s^{*}}=(s,s^{*})$. Roughly speaking, $\mathscr{R}_{0}(s,s^{*})$ can be interpreted as the expected number of secondary infections that result from any one infection. This can be computed via the next generation matrix method [19, 46, 47]: it is the spectral radius of the matrix $FV^{-1}$ where $F$ and $V$ are, respectively, the leading principal submatrices of $D\mathcal{F}(x_{s,s^{*}})$ and $D\mathcal{V}^{-}(x_{s,s^{*}})-D\mathcal{V}^{+}(x_{s,s^{*}})$ corresponding to the infected compartments, where $\mathcal{F},\mathcal{V}^{+},\mathcal{V}^{-}$ are as in (2.4),(2.5). In our case, these are
(2.10) $F=\begin{bmatrix}\beta(1-\alpha)s&\beta(1-\alpha)s\\ \beta s^{*}&\beta s^{*}\end{bmatrix},\hskip 56.9055ptV=\begin{bmatrix}\gamma+\eta+\delta+(\overline{\mu}-\mu)s^{*}&-\nu\\ -(\overline{\mu}-\mu)s^{*}&\gamma+\nu+\delta\end{bmatrix}$.
Because $F$ is rank 1, the spectral radius $\rho(FV^{-1})=\mathscr{R}_{0}(s,s^{*})$ can easily be computed by hand. We write it in two different equivalent ways to point out behavior with respect to changes in different parameters:
(2.11) $\begin{split}\mathscr{R}_{0}(s,s^{*})&=\frac{\beta\Big{[}(1-\alpha)\Big{(}\gamma+\nu+\delta+(\overline{\mu}-\mu)s^{*}\Big{)}s+\Big{(}\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*}\Big{)}s^{*}\Big{]}}{(\gamma+\delta)(\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*})+\eta\nu}\\ &=\frac{\beta\Big{[}(1-\alpha)\Big{(}1-\tfrac{\eta}{\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*}}\Big{)}s+s^{*}\Big{]}}{\gamma+\delta+\frac{\eta\nu}{\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*}}}.\end{split}$
The second representation lends itself to a reasonably natural interpretation, with the main complicating factor being the difference in recovery rates between the compliant and noncompliant populations due to the compliant population seeking treatment represented by $\eta$. In the worst case scenario, where $\xi=1,\nu=0$ so that the DFE is $(s,s^{*})=(0,\tfrac{b}{\delta})$, we find $\mathscr{R}_{0}(0,\tfrac{b}{\delta})=\frac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}.$ This is the reproductive ratio for the basic SIR model (2.1), which signifies that in this worst case scenario, the policy-maker’s control efforts do nothing to stop the spread of the disease. In the best case scenario, $\xi=0$ and $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\overline{\mu}-\mu}$ so that the only DFE is $(s,s^{*})=(\tfrac{b}{\delta},0)$ and we find that $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\frac{\beta(1-\alpha)}{\gamma+\eta+\delta}.$ This decreases the reproductive ratio when compared with the basic SIR model in two ways: $(1-\alpha)$ in the numerator results from the reduction in infectivity among the compliant susceptible population and the $\eta$ in the denominator results from the increased recovery rate among the compliant infected population. In general, $\mathscr{R}_{0}(s,s^{*})$ can then be viewed as a very complicated interpolation between these two endpoints which is illuminated a bit in the second line of (2.11) where the terms in front of $s$ in the numerator can be seen to represent reduction in disease spread among the compliant population, and that $s^{*}$ has no such modifications represents inefficacy of control efforts among the noncompliant population.
We make some natural observations regarding $\mathscr{R}_{0}(s,s^{*})$.
Proposition 2.2.
Define $\mathscr{R}_{0}(s,s^{*})$ as in (2.11) for nonnegative $s,s^{*}$.
  • (i)
    Holding other values constant, $\mathscr{R}_{0}(s,s^{*})$ is decreasing in each of the control variables $\alpha,\eta,\mu,\nu$.
  • (ii)
    The function $s^{*}\mapsto\mathscr{R}_{0}(\tfrac{b}{\delta}-s^{*},s^{*})$ is increasing for $s^{*}\in[0,\tfrac{b}{\delta}]$.
Remark. This proposition is a sanity check to verify that our model behaves as expected. The two points can be interpreted as follows: (i) more control leads to a lower rate of disease spread, and (ii) when we are at the equilibrium population level of $\tfrac{b}{\delta}$, more noncompliance leads to a higher rate of disease spread.
Proof.
That $\mathscr{R}_{0}(s,s^{*})$ is decreasing in $\alpha$ when other values are held constant is obvious directly from (2.11) since $\mathscr{R}_{0}(s,s^{*})$ is linear in $\alpha$. That $\mathscr{R}_{0}(s,s^{*})$ is decreasing in $\mu$ can be seen from the second line (2.11) since increasing $\mu$ increases each of the nested fractions, each of which has the effect of decreasing $\mathscr{R}_{0}(s,s^{*}).$ The remaining statements in (i), (ii) are not so obvious directly from the formula, but follow from the observation that $z\mapsto\frac{c_{1}+c_{2}z}{c_{3}+c_{4}z}$ is increasing for $z\geq 0$ when $c_{2}c_{3}\geq c_{1}c_{4}$ and decreasing for $z\geq 0$ when $c_{2}c_{3}\leq c_{1}c_{4}$. For (i), alternately writing $\mathscr{R}(s,s^{*})$ in this form where the role of $z$ is played by $\eta$ or $\nu$, the inequalities proving that $\mathscr{R}_{0}(s,s^{*})$ is decreasing boils down to
$(\gamma+\delta)(\gamma+\nu+\delta+(\overline{\mu}-\mu)s^{*})s^{*}$ $\leq(\gamma+\delta+\nu)(\gamma+\nu+\delta+(\overline{\mu}-\mu)s^{*})((1-\alpha)s+s^{*}),$
$(\gamma+\delta)(\gamma+\eta+\delta+(\overline{\mu}-\mu)s^{*})$ $\leq(\gamma+\eta+\delta)\Big{(}\gamma+\delta+(\overline{\mu}-\mu)s^{*}+\tfrac{\eta s^{*}}{(1-\alpha)s+s^{*}}\Big{)},$
respectively. Each of these is verified simply by checking that the right-hand side has all the same terms as the left hand side plus a few extra positive terms.
For part (ii), if $s^{*}\in[0,\tfrac{b}{\delta}]$, we see
(2.12) $\mathscr{R}_{0}(\tfrac{b}{\delta}-s^{*},s^{*})=\frac{\beta\Big{[}\tfrac{b}{\delta}(\gamma+\nu+\delta)+\Big{(}\tfrac{b}{\delta}(\overline{\mu}-\mu)+\eta\Big{)}s^{*}\Big{]}}{\Big{(}(\gamma+\delta)(\gamma+\eta+\nu+\delta)+\eta\nu\Big{)}+(\gamma+\delta)(\overline{\mu}-\mu)s^{*}}.$
That this is increasing in $s^{*}$ follows from the same reasoning with the inequality flipped. ∎
To conclude this section, we state stability results for our DFE. We state the theorem here, but relegate the proof to an appendix since it requires a long computation.
Theorem 2.3.
Considering the DFE values given in (2.8), (2.9) and $\mathscr{R}_{0}(s,s^{*})$ as defined in (2.11), we have the following.
  • (i)
    If $\xi=0$ and $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$, then $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$ and unstable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)>1$.
  • (ii)
    If $\xi=0$ and $\tfrac{b}{\delta}>\tfrac{\nu+\delta}{\overline{\mu}-\mu}$, then $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})<1$ and unstable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})>1$
  • (iii)
    If $\xi\in(0,1]$, then $(s_{3},s_{3}^{*})$ as defined in (2.9) is asymptotically stable when $\mathscr{R}_{0}(s_{3},s_{3}^{*})<1$ and unstable when $\mathscr{R}_{0}(s_{3},s_{3}^{*})>1$.
Remark. In each case, that the condition on the reproductive ratio determines local asymptotic stability or instability follows directly from [47, Theorem 2] once the assumptions of that theorem are verified. However, the stability is actually global, in the sense that under the conditions of each statement, any solution of (2.2) will tend toward the given equilibrium point regardless of whether the initial condition is nearby the equilibrium point. This global version can be proven using appropriately constructed Lyapunov functions as demonstrated in the appendix and in [39].
3. Optimal Control of our Model
[…]
With this notation, we define the marginal cost function,
(3.1) $r(x,u)=c_{1}(I+I^{*})+c_{2}N^{*}+\frac{c_{3}}{2}\alpha^{2}+\frac{c_{4}}{2}\eta^{2}+\frac{c_{5}}{2}\mu^{2}+\frac{c_{6}}{2}\nu^{2},$
where, as above, $N^{*}=S^{*}+I^{*}+R^{*}$. Our cost functional is then
(3.2) $J(x,u)=\int^{T}_{0}r(x(t),u(t))dt.$
In (3.1), $c_{1},\ldots,c_{6}$ are positive weights indicating the preferences of the policy-maker. For example, a more public-health oriented policy-maker would take $c_{1}$ larger, whereas an economically stingy policy-maker would increase $c_{3},c_{4},c_{5},c_{6}$. From the above, our admissible control set is
(3.3) $U_{\text{ad}}=\{(\alpha,\eta,\mu,\nu)\in{\mathbb{R}}^{4}\,:\,0\leq\alpha\leq 1,\,0\leq\eta\leq\overline{\eta},\,0\leq\mu\leq\overline{\mu},\,0\leq\nu\leq\overline{\nu}\}$
where $\overline{\eta},\overline{\mu},\overline{\nu}$ are fixed positive constants. We then consider control functions in
(3.4) $\mathscr{U}_{\text{ad}}=\{u(\cdot)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(\alpha(\cdot),\eta(\cdot),\mu(\cdot),\nu(\cdot)):[0,T]\to U_{\text{ad}}\,\,|\,\,u\text{ measurable}\}.$
Our optimal control problem is then
(3.5) $\min_{u\in\mathscr{U}_{\text{ad}}}J(x,u)\,\,\,\,\text{ subject to }\,\,\frac{dx}{dt}=f(x,u),\,\,\,\,x(0)=x_{0}$
where $f(x,u)$ is defined by the right-hand side of (2.2) and $x_{0}\in{\mathbb{R}}^{6}$ has nonnegative components whose sum is bounded by $\tfrac{b}{\delta}$.
Existence of an optimal control for this problem follows from classical optimal control theory.
Theorem 3.1.
Assuming $c_{3},c_{4},c_{5},c_{6}$ are positive and $x_{0}$ has nonnegative components, there exists an optimal control $u\in\mathscr{U}_{\text{ad}}$ which solves the minimization problem (3.5).
We omit the proof. Some key features of our system are that solutions of our system remain bounded and nonnegative as long as $x_{0}$ is componentwise nonnegative, $f(x,u)$ is smooth and affine in the control variables, and $r(x,u)$ is smooth and uniformly strongly convex in $u$ (when each of $c_{3},c_{4},c_{5},c_{6}$ are positive), and the control set $U_{\text{ad}}$ is convex and compact. These suffice to prove existence of an optimal control using [45, Theorem 2.1]. In fact the particular dependence of running cost (3.1) on the state variables $x=(S,I,R,S^{*},I^{*},R^{*})$ could be significantly altered, as long as the smoothness in all variables and uniform strong convexity in control variables are maintained. We choose this particular running cost only to represent the concerns we imagine a policy-maker may have; other decisions are certainly valid.
A necessary condition for optimality is provided by the Pontryagin minimum principle (PMP). Using state variables $x=(S,I,R,S^{*},I^{*},R^{*})$, costate variables $p=(p_{S},p_{I},p_{R},p_{S^{*}},p_{I^{*}},p_{R^{*}})$, and control variables $u=(\alpha,\eta,\mu,\nu),$ we define the Hamiltonian
(3.6) $\begin{split}H(x,p,u)&=\langle p,f(x,u)\rangle+r(x,u),\end{split}$
where $r(x,u)$ is the running cost (3.1) and $f(x,u)$ is defined by the right-hand side of (2.2). Then the PMP states that if $x(\cdot),u(\cdot)$ are an optimal state and control pair for (3.5), then there is a corresponding costate trajectory $p(\cdot)$ such that
(3.7) $\frac{dx}{dt}$ $=\hphantom{-}\nabla_{p}\mathcal{H}(x,p,u),\,\,\,\,\,\,\,\,\,\,x(0)=x_{0},$
(3.8) $\frac{dp}{dt}$ $=-\nabla_{x}\mathcal{H}(x,p,u),\,\,\,\,\,\,\,\,\,\,p(T)=0,$
(3.9) $u(t)$ $=\operatorname*{argmin}_{\hat{u}\in U_{\text{ad}}}\mathcal{H}(x(t),p(t),\hat{u}),\,\,\,\,\,\text{ for all }\,\,\,t\in[0,T].$
We direct the reader to several classical sources for proofs and discussions regarding the PMP; for example, [2, §5], [29, §4], [42, §1,2], [49, §12]. We note that (3.7) is exactly the original dynamics (2.2). We refer to (3.8) as the costate dynamics, and (3.9) as the optimality condition. The condition $p(T)=0$ in (3.8) is called the transversality condition (discussed in the aforementioned sources), and takes this simplified form because we consider a fixed time, free end-point problem with no terminal cost.
We will use the forward-backward sweep method described in [28, §4] to numerically solve (3.7)-(3.9). We fully detail the algorithm shortly, but first, to make all computations entirely explicit, we write out (3.8) and (3.9) for our specific choice of cost functional (3.2). In our case, the Hamiltonian can be written
(3.10) $\begin{split}\mathcal{H}(x,p,u)&=p_{S}\Big{(}(1-\xi)b-\beta(1-\alpha)S(I+I^{*})-(\overline{\mu}-\mu)SN^{*}+\nu S^{*}-\delta S\Big{)}\\ &\hskip 28.45274pt+p_{I}\Big{(}\beta(1-\alpha)S(I+I^{*})-(\gamma+\eta)I-(\overline{\mu}-\mu)IN^{*}+\nu I^{*}-\delta I\Big{)}\\ &\hskip 28.45274pt+p_{R}\Big{(}(\eta+\gamma)I-(\overline{\mu}-\mu)RN^{*}+\nu R^{*}-\delta R\Big{)}\\ &\hskip 28.45274pt+p_{S^{*}}\Big{(}\xi b-\beta S^{*}(I+I^{*})+(\overline{\mu}-\mu)SN^{*}-\nu S^{*}-\delta S^{*}\Big{)}\\ &\hskip 28.45274pt+p_{I^{*}}\Big{(}\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu)IN^{*}-\nu I^{*}-\delta I^{*}\Big{)}\\ &\hskip 28.45274pt+p_{R^{*}}\Big{(}\gamma I^{*}+(\overline{\mu}-\mu)RN^{*}-\nu R^{*}-\delta R^{*}\Big{)}\\ &\hskip 28.45274pt+c_{1}(I+I^{*})+c_{2}N^{*}+c_{3}\frac{\alpha^{2}}{2}+c_{4}\frac{\eta^{2}}{2}+c_{5}\frac{\mu^{2}}{2}+c_{6}\frac{\nu^{2}}{2}.\end{split}$
Thus, (3.8) reads
(3.11) $\begin{split}\frac{dp_{S}}{dt}&=\beta(1-\alpha)(I+I^{*})(p_{S}-p_{I})+(\overline{\mu}-\mu)N^{*}(p_{S}-p_{S^{*}})+\delta p_{S},\\ \frac{dp_{I}}{dt}&=\beta(1-\alpha)S(p_{S}-p_{I})+\beta S^{*}(p_{S^{*}}-p_{I^{*}})+(\gamma+\eta)(p_{I}-p_{R})+\delta p_{I}-c_{1}\\ &\hskip 28.45274pt+(\overline{\mu}-\mu)N^{*}(p_{I}-p_{I^{*}}),\\ \frac{dp_{R}}{dt}&=(\overline{\mu}-\mu)N^{*}(p_{R}-p_{R^{*}})+\delta p_{R},\\ \frac{dp_{S^{*}}}{dt}&=\beta(I+I^{*})(p_{S^{*}}-p_{I^{*}})+\nu(p_{S^{*}}-p_{S})+\delta p_{S^{*}}-c_{2}\\ &\hskip 28.45274pt+(\overline{\mu}-\mu)\big{(}S(p_{S}-p_{S^{*}})+I(p_{S}-p_{S^{*}})+R(p_{R}-p_{R^{*}})\big{)},\\ \frac{dp_{I^{*}}}{dt}&=\beta(1-\alpha)S(p_{S}-p_{I})+\beta S^{*}(p_{S^{*}}-p_{I^{*}})+\gamma(p_{I^{*}}-p_{R^{*}})+\nu(p_{I^{*}}-p_{I})-c_{1}-c_{2}\\ &\hskip 28.45274pt+(\overline{\mu}-\mu)\big{(}S(p_{S}-p_{S^{*}})+I(p_{S}-p_{S^{*}})+R(p_{R}-p_{R^{*}})\big{)}+\delta p_{I^{*}},\\ \frac{dp_{R^{*}}}{dt}&=\nu(p_{R^{*}}-p_{R})+(\overline{\mu}-\mu)\big{(}S(p_{S}-p_{S^{*}})+I(p_{S}-p_{S^{*}})+R(p_{R}-p_{R^{*}})\big{)}+\delta p_{R^{*}}-c_{2},\end{split}$
along with zero terminal condition $p_{X}(T)=0$ for each $X\in\{S,I,R,S^{*},I^{*},R^{*}\}.$
Because the Hamiltonian is quadratic in the control variables $u=(\alpha,\eta,\mu,\nu)$, one can easily resolve the optimality condition (3.9) by simply minimizing a quadratic function and projecting back onto the admissible set of control values. Suppressing dependence on $t$, once the state and costate variables are known, optimal controls are given by
(3.12) $\begin{split}\alpha&=\min\left\{\max\left\{\frac{\beta S(I+I^{*})(p_{I}-p_{S})}{c_{3}},\,0\right\},\,1\right\},\\ \eta&=\min\left\{\max\left\{\frac{I(p_{I}-p_{R})}{c_{4}},\,0\right\},\,\overline{\eta}\right\},\\ \mu&=\min\left\{\max\left\{\frac{[S(p_{S^{*}}-p_{S})+I(p_{I^{*}}-p_{I})+R(p_{R^{*}}-p_{R})]N^{*}}{c_{5}},\,0\right\},\,\overline{\mu}\right\},\\ \nu&=\min\left\{\max\left\{\frac{S^{*}(p_{S^{*}}-p_{S})+I^{*}(p_{I^{*}}-p_{I})+R^{*}(p_{R^{*}}-p_{R})}{c_{6}},\,0\right\},\,\overline{\nu}\right\}.\end{split}$
Given this, the forward-backward sweep method is detailed in algorithm 1. We include some basic implementation notes. For the initial guess, we always use $u^{(0)}\equiv 0$. Step (iv) in the iteration is a relaxation step which may not be necessary but, empirically, accelerates convergence. For convergence criterion, we halt the iteration when the relative change in control variables (measured in supremum norm) is below some specified tolerance. Using the suggestion of [28, §4], we rearrange the condition slightly so as to avoid division by zero in the case that zero control is optimal. Thus our convergence criterion is
(3.13) $(\text{tol})\cdot\|\alpha^{(k+1)}\|_{\infty}-\|\alpha^{(k+1)}-\alpha^{(k)}\|_{\infty}\geq 0,$
where “tol” is the convergence tolerance. While (3.13) is written for $\alpha$, we enforce this criterion for each of $\alpha,\eta,\mu,\nu$ and only halt the iteration once all of them satisfy the condition.
Algorithm 1 Forward-Backward Sweep Method for Resolving Optimal Controls
Input: initial condition $x_{0}=(S_{0},I_{0},R_{0},S^{*}_{0},I^{*}_{0},R^{*}_{0})\in{\mathbb{R}}^{6}_{+}$, parameter values $b,\delta,\beta,\gamma,\overline{\eta},\xi,\overline{\mu},\overline{\nu}$, cost weights $c_{1},\ldots,c_{6}$, a relaxation parameter $\kappa\in[0,1]$ and an initial guess for the control map $u^{(0)}=(\alpha^{(0)},\eta^{(0)},\mu^{(0)},\nu^{(0)})$. Set $k=0$ and do the following:
repeat
  (i) Resolve $x^{(k+1)}=(S^{(k+1)},I^{(k+1)},R^{(k+1)},(S^{*})^{(k+1)},(I^{*})^{(k+1)},(R^{*})^{(k+1)})$ as the solution of (2.2) with initial condition $x_{0}$, using control variables $u^{(k)}$
  (ii) Resolve $p^{(k+1)}=(p_{S}^{(k+1)},p_{I}^{(k+1)},p_{R}^{(k+1)},p_{S^{*}}^{(k+1)},p_{I^{*}}^{(k+1)},p_{R^{*}}^{(k+1)})$ as the solution of (3.11) with zero terminal condition, using control variables $u^{(k)}$ and state variables $x^{(k+1)}$
  (iii) Define $u^{*}=(\alpha^{(k+1)},\eta^{(k+1)},\mu^{(k+1)},\nu^{(k+1)})$ from (3.12) using state variables $x^{(k+1)}$ and costate variables $p^{(k+1)}$
  (iv) Define $u^{(k+1)}=(1-\kappa)u^{(k)}+\kappa u^{*}$ and increment $k$
until convergence criteria are met
return $u=u^{(k)}$, $x=x^{(k)}$, $p=p^{(k)}$: respectively the optimal control, state, and costate trajectories
4. Results & Discussion
We implement algorithm 1 in MATLAB and demonstrate the results in several different cases. Specifically, we are interested in observing the differing behavior of optimal control maps in response to changes in the policy-maker’s preferences, expressed via the weights $c_{1},\ldots,c_{6}$ from (3.1). For all simulations, we use relaxation parameter $\kappa=0.8$ and a time horizon of $T=100$, and for the sake of numerically resolving the solutions of (2.2) and (3.11), we use an explicit Euler scheme with uniform discretization of $[0,T]$ with step size $\Delta t=0.1$, and a convergence tolerance of $\text{tol}=10^{-3}.$ The code which generated all of these results is available online. [^1] https://github.com/chparkinson/Opt_Control_SIR_With_Noncompliance_As_Social_Contagion
What remains is to specify initial conditions and the parameters $b,\delta,\beta,\gamma,\overline{\eta},\xi,\overline{\mu},\overline{\nu}$. To enhance readability, we include a description of each of these parameters and the values they take in each scenario. In all scenarios, we take $b=\delta$, and initial condition $(S_{0},I_{0},R_{0},S^{*}_{0},I^{*}_{0},R^{*}_{0})=(0.69,0.01,0,0.29,0.01,0)$, indicating that at the outset of the epidemic, 2% of the population is infected, and 30% of the population is noncompliant. Because $b=\delta$ and the initial condition has total population $1$, by (2.6), the total population will be $1$ for all time, so individual compartments can be interpreted as percentages of the total population.
Before displaying results, we briefly discuss the bearing of the controls. Recall that the cost (3.2) for the policy-maker is split into three parts: (C1) cost associated with infections (weighted by the constant $c_{1}$), (C2) costs associated with noncompliance (weighted by the constant $c_{2}$) and (C3) cost associated with controls (weighted by constants $c_{3},\ldots,c_{6}$). By proposition 2.2, any use of controls will help decrease cost (C1) by slowing the spread of infections. This is the only effect that $\alpha$ and $\eta$ can have. The controls $\mu$ and $\nu$ obliquely lower infection rate and thus reduce cost (C1), but more explicitly decrease cost (C2) by decreasing the noncompliant population. Of course, any use of control increases cost (C3), meaning that the policy-maker has a trade-off to consider. We mention specifically that the absolute size of the weights is not important, but rather their size relative to each other. Accordingly, in much of the below we fix the control cost weights $c_{3},\ldots,c_{6}=1$ and vary the cost weights $c_{1},c_{2}$ associated with total infections and total amount of noncompliance, respectively. In In certain cases, we would like to “turn off” individual controls; to do so, one can simply set that control to zero (this could equivalently be accomplished by sending the control’s cost weight to infinity). Lastly, as baseline values for infection and noncompliance cost weights, we choose $c_{1}=1$ and $c_{2}=0.1$. This causes the cost due to infection and the cost due to noncompliance to be roughly equal in the uncontrolled case. The reason that $c_{2}$ must be significantly smaller is that with no controls, noncompliance is a permanent state, whereas infection is not, which means that noncompliance has the potential to contribute much more to the total cost.
Parameter Description Scen. 1 Scen. 2 Scen. 3
$b$ natural birth rate 0.01 0.01 0.01
$\delta$ natural death rate 0.01 0.01 0.01
$\beta$ disease infection rate 0.4 0.6 0.4
$\gamma$ disease recovery rate 0.2 0.2 0.2
$\overline{\eta}$ maximum achievable treatment efficacy 0.1 0.1 0.1
$\xi$ noncompliant portion of new population 0 0.3 var.
$\overline{\mu}$ baseline infection rate for noncompliance 0.1 0.2 0.5
$\overline{\nu}$ maximum achievable rehab. rate for noncompliance 0.1 0.1 0.1
$c_{1}$ cost weight on infections var. var. 1
$c_{2}$ cost weight on infections var. var. 0.1
$c_{3},\ldots,c_{6}$ cost weights on controls 1 1 1
Figure 1. Table of parameters for each scenario. Recall in particular that $\overline{\eta},\overline{\mu},\overline{\nu}$ are upper bounds for the control variables $\eta,\mu,\nu$, which represent increase in recovery rate due to seeking treatment, decrease in spread rate of noncompliance due to public health initiatives, and recovery rate for noncompliance due to educational campaigns, respectively. Not mentioned is the control variable $\alpha$, representing decrease in infection rate among the compliant susceptible population due to nonpharmaceutical intervention measures, since $\alpha$ will take values in $[0,1].$
4.1. Scenario 1
The parameters for this scenario are listed in figure 1, and the result of the simulation in the absence of controls is included in figure 2. With no controls, the reproductive ratio for this scenario is $\mathscr{R}_{0}=\tfrac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}=\frac{0.4}{0.21}\approx 1.9.$ We note in particular that $\xi=0$, meaning that all newly introduced members of the population are compliant, and we are in the regime of theorem 2.3(i),(ii). In this case, there are two DFEs, that, when controls are constant, take the forms $(s_{1},s^{*}_{1})=(\tfrac{b}{\delta},0)$ and $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$. If the controls are constant, the second is stable in the case that it is physically meaningful (i.e. $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$); otherwise, the first is stable. Clearly to minimize (C1) and (C2), it behooves the policy-maker to drive the system toward $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$. If control cost is no issue, the policy-maker can increase $\mu,\nu$ to ensure that $(s_{1},s_{1}^{*})$ is stable, and then decrease $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\tfrac{\beta(1-\alpha)}{\gamma+\eta+\delta}$ using $\alpha,\mu$ so as to fall into the stability regime of theorem 2.3(i). However, it will likely be too expensive in terms of (C3) to actually fully do this. In figure 3, we see the optimal control plans for a “balanced” policymaker: one who would like to minimize both infections and noncompliance. The disease dynamics are plotted in the top panels figure 3, split into compliant and noncompliant compartments. The bottom left panel of figure 3 displays the optimal control plans. Finally, the bottom right displays the reproductive ratio $\mathscr{R}_{0}(s,s^{*})$ which, due to the controls, now changes with time. In particular, we plot the value of $\mathscr{R}_{0}(\tfrac{b}{\delta},0)$ when $\tfrac{b}{\delta}<\frac{\nu+\delta}{\overline{\mu}-\mu}$ to reflect the conditions of theorem 2.3(i) and $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$ when $\tfrac{b}{\delta}\geq\frac{\nu+\delta}{\overline{\mu}-\mu}$ to reflect the conditions of theorem 2.3(ii). In either case, if the corresponding reproductive ratio was kept below 1, the corresponding DFE would be stable, but from the plot we see that it is not optimal to do so. Instead, controls $\mu$ and $\nu$ are used to ensure compliance in the early stages of the epidemic, so that $\alpha$ and $\eta$ can be effectively deployed to keep the reproductive ratio low, though a small outbreak is indeed allowed to occur. With these cost weights, the optimal controls reduce the total cost from $12.42$ in the uncontrolled case to $2.97$ in the optimally controlled case, representing a roughly $76\%$ relative reduction in cost due to controls.
Figure 2. Simulation of Scenario 1 in the absence of controls.
Figure 3. Simulation of Scenario 1 with a “balanced” policy-maker: cost weights $c_{1}=1,c_{2}=0.1$. The strategy is to quell noncompliance using $\mu,\nu$ so as to keep the infection rate low. Since the majority of the population is compliant, $\alpha$ and $\nu$ can be deployed to effectively lower the reproductive ratio. However, it is not lowered so much that $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1.$ A small outbreak of the disease is allowed to occur.
Figure 4. Simulation of Scenario 1 with a public health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0$. In this case, even though there is no explicit cost associated with noncompliance, because the policy-maker desires to keep infections low, the optimal strategy employs $\mu$ and $\nu$ even more strongly than in the balanced case. The disease controls $\alpha$ and $\eta$ are also increased so as to drive $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$ and take advantage of the stability of the DFE $(\tfrac{b}{\delta},0)$ until the disease has been effectively eradicated.
We contrast this with the optimal actions corresponding to cost weights $c_{1}=5,c_{2}=0$. This represents a public health oriented policy-maker whose primary concern is to stop the disease, and does not even assess cost of noncompliance. The results of the simulation using these weights are in figure 4. In this case, the cost weight on infections is high enough that it is indeed optimal to use controls $\mu,\nu$ to ensure that, in the initial stages of the epidemic, we remain in the stability regime for the disease-free equilibrium $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$, and then use $\alpha,\eta$ to drive $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$, in essense taking advantage of theorem 2.3(i). This is maintained until the disease has been effectively eradicated, at which time the policy-maker eases the controls (note that, in the context of this model, the policy-maker is aware of the horizon time $T=100$, so $\alpha$ is tapered off such that infections will not significantly increase within this final time). Figure 4 demonstrates the need for synergy between the disease controls $\alpha,\eta$ and noncompliance controls $\mu,\nu$. With these cost weights, there is no explicit cost associated with increased noncompliance. However, because increased noncompliance accelerates disease spread, the policy-maker must still employ $\mu,\nu$ to keep noncompliance low so as to accomplish their goal of disease elimination. With these cost weights, the total cost is reduced from $22.60$ in the uncontrolled case to $6.90$ in the optimally controlled case, representing a roughly $70\%$ relative reduction in cost.
To further explore the synergy between the controls, we include one final figure using the scenario 1 parameters, and the health oriented policy-maker with cost weights $c_{1}=5,c_{2}=0$. In this case, we demonstrate the effects of “turning off” certain controls. The results are in figure 5, where we display the effects on the optimal control maps of forcing $\eta=0$ (top), or forcing $\mu=\nu=0$ (bottom). Recall, $\eta$ represents treatment efforts that hasten recovery among the compliant infected population, so one could think of $\eta=0$ as a scenario near the outset of an epidemic where treatment is unavailable or ineffective. In this case, the policy-maker’s optimal strategy is to more strongly enforce NPIs (increase $\alpha$), while also using $\mu,\nu$ to ensure compliance. While they are not displayed, the dynamics in this case look, in essence, the same as those in the top panel of figure 4 in that the disease is effectively eradicated. However, without access to the control $\eta$, the total cost can only be reduced from $22.60$ in the uncontrolled case to $15.43$ in the optimally controlled case, representing only a roughly $32\%$ relative cost reduction (compared to a $70\%$ relative cost reduction when $\eta$ is available). Thus, in some sense, controls efforts are only half as effective without $\eta$.
Figure 5. Simulation of Scenario 1 with a health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0$, with either treatment $\eta$ unavailable (top), or control of noncompliance $\mu,\nu$ unavailable (bottom). Note the $y$-axis change in the top left.
In a similar vein, $\mu=\nu=0$ represents a scenario wherein it is impossible to encourage compliance or slow the spread of noncompliance. In this case, the dynamics look in essence like those in figure 2 where no controls are used. One sees in the bottom panel of figure 5 that without the ability to curb noncompliance, it is still beneficial to slightly decrease the reproductive ratio among the compliant population using $\alpha$ and $\eta$, but much less so since the majority of the population will become noncompliant, hampering disease prevention efforts. With $\mu=\nu=0$, one cannot land in the stability regime of theorem 2.3(i), nor accomplish $\mathscr{R}_{0}(\frac{\nu+\delta}{\overline{\mu}-\mu},\frac{b}{\delta}-\frac{\nu+\delta}{\overline{\mu}-\mu})<1$, so the disease is simply allowed to run its course. The use of $\alpha$ and $\eta$ still accomplish a reduction of total cost from $22.60$ in the uncontrolled case to $19.69$ in the optimally controlled case, representing a roughly $13\%$ relating reduction in total cost (compared to the $70\%$ cost reduction when all controls are available). Meaning that, with $\mu=\nu=0$, controls efforts are only about one-fifth as effective.
4.2. Scenario 2
In this scenario, when compared with scenario 1, we increase the disease infection rate $\beta$ and the spread rate $\overline{\mu}$ of noncompliance, and we also assume that $30\%$ of the newly introduced population members are noncompliant ($\xi=0.3$), while all other parameters are unchanged. For this scenario, we set $c_{2}=0.01$, indicating a weak preference of the policy-maker to quell noncompliance for its own sake, though quelling noncompliance for the sake of slowing disease spread is still desirable. In figure 6, we display the results of the simulation with $c_{1}$ successively taking values $1/3,1,3,6,9$. In this case, noncompliance is much more difficult to control since 30% of the newly introduced population is noncompliant, and noncompliance spreads fairly quickly. When noncompliance is difficult to control, the control variables $\alpha$ and $\eta$ have a diminished effect. Accordingly, in this parameter regime, even a strongly public-health oriented policy-maker will not fully suppress disease spread. Note, for example, that the dynamics are qualitatively similar in the cases of $c_{1}=1$ and $c_{1}=6$. It isn’t until $c_{1}=9$ that the policy-maker significantly increases the use of the control $\alpha$ so as to more effectively stunt disease spread. This indicates that in scenarios where disease spread is fast enough, and noncompliance is rampant enough, a policy-maker who is even slightly economically minded (i.e. concerned with keeping control costs down) will let the disease run its course.
Figure 6. Simulation of Scenario 2, with increased spread of the disease an noncompliance. Noncompliance cost weight is $c_{2}=0.01$. From top to bottom, the cost weight on infections is $c_{1}=1/3,1,3,6,9$. Because disease spread is so fast, suppression efforts are costly and it is not optimal to fully suppress the disease until $c_{1}$ is near 99. In a case like this, an economically oriented policy-maker would allow the disease to spread.
4.3. Scenario 3
Finally, in this scenario, we revert to the parameters of the baseline simulation in figure 3, except that we increase the spread rate for noncompliance to $\overline{\mu}=0.5$ and observe the effects of changing $\xi$, the proportion of the newly introduced population that is noncompliant. In particular, we have the “balanced” cost weights $c_{1}=1,c_{2}=0.1$. The full parameter list for scenario 3 is in figure 1. In some sense, $\xi$ could be interpreted as the baseline proclivity toward noncompliance within the population. A key question is then the level at which this proclivity is high enough that it is no longer optimal to attempt to suppress noncompliance. We include four plots in figure 7 where $\xi$ takes successive values $0,1/4,1/2,1.$ When $\xi=0$, even with the very high rate of noncompliance spread, it is optimal to fully quell noncompliance in the early stages of the epidemic. This is seen, for example, in the top panels of figure 7, where $\mu$ still takes its maximum value for a short time at the beginning, and noncompliance stays throughout most of the time interval (we note again that it only grows at the end because the policy-maker is aware of the horizon time, and can save some cost by releasing some controls when there is no longer time for an outbreak to occur). This is contrasted with the ensuing plots when $\xi=1/4,1/2,1$. In those cases, $\mu$ never takes it maximum value, and when $\xi=1/2,1$, $\mu$ remains below half its maximum value. The results for the cases $\xi=1/2$ and $\xi=1$ are very similar, seeming to indicate that for values $\xi\geq 1/2$, controls are simply ineffective. We note also that $\alpha$ and $\nu$ decrease with increasing $\xi$, which is a result of the fact that disease control is less effective when control of noncompliance cannot be achieved.
Figure 7. Simulation of Scenario 3, with increased spread rate of noncompliance, balanced cost weights, and variable proportion of the newly introduced population which is noncompliant $(\xi)$. Even with the increased spread of noncompliance, as long as $\xi=0$, the optimal strategy still employs the maximum available $\mu$ for a small time to suppress spread of noncompliance. As $\xi$ increases, slowing spread of noncompliance becomes futile, so that even when $\xi=0.25$ the policy-maker no longer uses the maximum value of $\mu$, and when $\xi=1/2$, $\mu$ is never more than half its maximum value and essentially the entire epidemic occurs on the noncompliant side. There is very little change between the simulations using $\xi=1/2$ and $\xi=1$, indicating that by the time $\xi\geq 1/2$, a threshold has been crossed and controls are ineffective.
5. Conclusion & Future Directions
In this paper, we present and analyze a compartmental SIR-style model for epidemiology, incorporating noncompliance with governmental NPIs as a social contagion. We provide stability analysis for the disease-free equilibria of the model, and then introduce four control variables that a policy-maker may have access to: (1) reduction in infectivity corresponding to the strength of the NPIs the policy-makers enacts, (2) increase in recovery rate for compliant individuals corresponding to increased treatment efforts, (3) reduction in spread of the social contagion of noncompliance corresponding to public health information campaigns, and (4) “recovery” from noncompliance corresponding to educational campaigns aimed at increasing compliance. We suggest a cost functional which is quadratic in controls and jointly linear in total number of infections and total amount of noncompliance. We give a characterization of the optimal controls via the Pontryagin optimality principle. Finally, we demonstrate the behavior of our model, focusing specifically on the behavior of the optimal control maps in several parameter regimes that represent different physical scenarios and policy-maker preferences.
We suggest several avenues for future work. A first would be to incorporate stochastic effects to account for various uncertainties inherent to both epidemic spread of diseases and human behavior. This is partially addressed in [39], but not in the context of optimal control, and quantification of the effect that stochasticity has on the optimal control maps could be interesting. Another direction would be to synthesize these results with other modeling frameworks. For example, a multiplex network describing the coevolution of disease and opinion spread is presented in [41], but does not include optimal control of disease spread. Modeling controls in a similar manner to ours but in the network setting could help illuminate the ways in which heterogeneity of contact networks affects spread of the disease in the presence of noncompliant behavior. Finally, inclusion of more finely modeled control variables could push models like this toward real-world utility. This could include things like the singular and/or one-shot vaccinations as in [17, 26] respectively, limited quarantine availability in [3], or control of information via media outlets like in [18]. Incorporation of finely tuned controllers into models of epidemiology with human behavioral effects could lead to vastly increased modeling fidelity.
Acknowledgments
CN was partially supported by the Honors Research Assistant Program (HRAP) of the University of Oklahoma. WW was partially supported by the Simons Foundation grant. Both CN and WW were partially supported by the Junior Faculty Fellowship from the University of Oklahoma.
Appendix A Proof of Theorem 2.3
In this appendix, we prove theorem 2.3. As stated in the remark below theorem 2.3, if one is willing to settle for local asymptotic stability in the case that $\mathscr{R}_{0}(s,s^{*})<1$, the theorem follows directly from [47, Theorem 2], whose hypotheses we recall and verify shortly. One can then upgrade to global stability using Lyapunov functions. We include both arguments. Recall, for the sake of this theorem, we are treating all control variables as constant. We also take $\mu\in[0,\overline{\mu})$. In the case the $\mu=\overline{\mu}$, the same results hold under the convention that $\overline{\mu}-\mu\sim 0^{+}$, so that $\tfrac{1}{\overline{\mu}-\mu}\sim+\infty$. In this case, the hypotheses of theorem 2.3(ii) cannot be met.
Proof.
Recall that we consider the variables $x=(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=(I,I^{*},S,R,S^{*},R^{*})$ and the infection and transfer functions $\mathcal{F},\mathcal{V}^{+},\mathcal{V}^{-}$ defined in (2.4), (2.5).
Defining $\mathscr{R}_{0}$ as the spectral radius of the matrix $FV^{-1}$ as described above, Theorem 2 of [47] gives local asymptotic stability of a DFE for a compartmental epidemic model when $\mathscr{R}_{0}<1$ and instability when $\mathscr{R}_{0}>1$ as long as the following hypotheses are satisfied:
  • (H1)
    $\mathcal{F}(x),\mathcal{V}^{+}(x),\mathcal{V}^{-}(x)\geq 0$ whenever $x$ is componentwise nonnegative.
  • (H2)
    For $i=1,2$, if $x_{i}=0$, then $\mathcal{V}_{i}^{-}(x)=0$.
  • (H3)
    For $i=3,4,5,6$, $\mathcal{F}_{i}(x)\equiv 0$.
  • (H4)
    If $x_{1}=x_{2}=0$, then $\mathcal{F}_{1}(x)=\mathcal{F}_{2}(x)=\mathcal{V}^{+}_{1}(x)=\mathcal{V}^{+}_{2}(x)=0.$
  • (H5)
    At the DFE $x$, all eigenvalues of $D\mathcal{V}^{-}(x)-D\mathcal{V}^{+}(x)$ have positive real part.
Note that (H1)-(H4) are general hypotheses that the system must satisfy (and that our system is easily seen to satisfy), whereas (H5) is a condition specific to the DFE that one is analyzing. Toward verifying (H5), we see that at a DFE $x=(0,0,s,0,s^{*},0)$, $M\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=D\mathcal{V}^{-}(x)-D\mathcal{V}^{+}(x)$ has the form
(A.1) $M=\left[\begin{smallmatrix}&&&&&\\ \gamma+\eta+\delta+(\overline{\mu}-\mu)s^{*}&-\nu&0&0&0&0\\ -(\overline{\mu}-\mu)s^{*}&\gamma+\nu+\delta&0&0&0&0\\ \beta(1-\alpha)s&(\beta(1-\alpha)+(\overline{\mu}-\mu))s&\delta+(\overline{\mu}-\mu)s^{*}&0&(\overline{\mu}-\mu)s-\nu&(\overline{\mu}-\mu)s\\ -(\gamma+\eta)&0&0&\delta+(\overline{\mu}-\mu)s^{*}&0&-\nu\\ \beta s^{*}&\beta s^{*}-(\overline{\mu}-\mu)s&-(\overline{\mu}-\mu)s^{*}&0&\nu+\delta-(\overline{\mu}-\mu)s&-(\overline{\mu}-\mu)s\\ 0&-\gamma&0&-(\overline{\mu}-\mu)s^{*}&0&\nu+\delta\\ &&&&&\end{smallmatrix}\right]$
To find all six eigenvalues of this matrix, we note that the two eigenvalues of the first principal $2\times 2$ submatrix will also be eigenvalues of $M^{T}$ with eigenvectors of the form $v=(v_{1},v_{2},0,0,0,0)$, and thus they are also eigenvalues of $M$. Since $\gamma+\delta>0$, these eigenvalues are necessarily positive; this follows from Gershgorin’s theorem since the first principal $2\times 2$ submatrix is strictly (columnwise) diagonally dominant.
The other four eigenvalues of $M$ have eigenvectors of the form $v=(0,0,v_{3},v_{4},v_{5},v_{6})$ and can be found by zooming into the last principal $4\times 4$ submatrix. Computing by hand, we find that these eigenvalues are
$\lambda_{1}=\delta,\,\,\,\,\,\,\,\,\,\,\lambda_{2}=\delta,\,\,\,\,\,\,\,\,\,\,\lambda_{3}=\delta+\nu+(\overline{\mu}-\mu)s^{*},\,\,\,\,\,\,\,\,\,\,\lambda_{4}=\delta+\nu+(\overline{\mu}-\mu)(s^{*}-s).$
These are all positive except possibly $\lambda_{4}$ so (H5) simply requires that $\lambda_{4}>0$, or
(A.2) $s-s^{*}<\frac{\nu+\delta}{\overline{\mu}-\mu}.$
When $\xi=0$, for $(s_{1},s^{*}_{1})=(\tfrac{b}{\delta},0)$, (A.2) reduces to $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$ which is the additional hypothesis of statement (i).
When $\xi=0$, for $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$, (A.2) reduces to $\tfrac{b}{\delta}>\tfrac{\nu+\delta}{\overline{\mu}-\mu},$ which is the additional hypothesis of statement (ii).
When $\xi\in(0,1]$, for $(s_{3},s_{3}^{*})$ as defined in (2.9), (A.2) reduces to
$\tfrac{\nu+\delta}{\overline{\mu}-\mu}-\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu}\right)^{2}+\tfrac{4\xi b}{\overline{\mu}-\mu}}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$
which is always satisfied, meaning no additional hypotheses are needed.
Thus (i), (ii), (iii) follow from [47, Theorem 2] if the statements are weakened to read “local asymptotic stability.”
Upgrading this to global asymptotic stability is more difficult, but can be accomplished using suitably constructed Lyapunov functions as in [39], where a stochastic model similar to (2.2) is analyzed. So as to keep this manuscript mostly self-contained, but also refrain from repeating a very long computation, we give the construction of an explicit Lyapunov function for $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$ in the case that $\xi=0$ and $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$, and refer the reader to [39] for other cases. The other cases are similar, but in some cases, the construction of the Lyapunov function relies on [47, lemma 6], which says, in essence, that under assumptions (H1)-(H5), $\mathscr{R}_{0}=\rho(FV^{-1})<1$ if and only if all eigenvalues of $M\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=F-V$ have negative real part where $F,V$ are as in (2.10). In turn, by the theory of Sylvester equations, this implies the existence of a unique symmetric, positive definite matrix $Q$ such that $M^{t}Q+QM=-I$, and the Lyapunov function is constructed using this matrix $Q$. The details are in [39].
For an explicit construction, suppose $\xi=0$ and $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$ so that the relevant DFE is $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$. Let $(S,I,R,S^{*},I^{*},R^{*})$ be any solution of (2.2) and set $\Sigma=S-\tfrac{b}{\delta}$. Recall, solution of (2.2) are nonnegative and bounded by $\tfrac{b}{\delta}$, so $\Sigma\leq 0$ and $I,R,S^{*},I^{*},R^{*}\geq 0$. Writing (2.2) with this new variable $\Sigma$ (and setting $\xi=0$), we see
(A.3) $\begin{split}\tfrac{d\Sigma}{dt}&=-\beta(1-\alpha)(\Sigma+\tfrac{b}{\delta})(I+I^{*})-(\overline{\mu}-\mu)(\Sigma+\tfrac{b}{\delta})N^{*}+\nu S^{*}-\delta u,\\ \tfrac{dI}{dt}&=\beta(1-\alpha)(\Sigma+\tfrac{b}{\delta})(I+I^{*})-(\gamma+\eta)I-(\overline{\mu}-\mu)IN^{*}+\nu I^{*}-\delta I,\\ \tfrac{dR}{dt}&=(\gamma+\eta)I-(\overline{\mu}-\mu)RN^{*}+\nu R^{*}-\delta R,\\ \tfrac{dS^{*}}{dt}&=-\beta S^{*}(I+I^{*})+(\overline{\mu}-\mu)(\Sigma+\tfrac{b}{\delta})N^{*}-\nu S^{*}-\delta S^{*},\\ \tfrac{dI^{*}}{dt}&=\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu)IN^{*}-\nu I^{*}-\delta I^{*},\\ \tfrac{dR^{*}}{dt}&=\gamma I^{*}+(\overline{\mu}-\mu)RN^{*}-\nu R^{*}-\delta R^{*}.\end{split}$
We want to prove that the trivial solution of (A.3) is globally stable, so we define a positive definite, radially coercive Lyapunov function $V(y)$, where $y=(\Sigma,I,R,S^{*},I^{*},R^{*})$ such that $\frac{d}{dt}V(y(t))\leq-cV(y(t))$ for some positive constant $c$. The particular form of $V(y)$ is certainly non-unique. So as to cancel some nonlinear terms in the computations, we define
(A.4) $V(y)=(\Sigma+I+S^{*}+I^{*})^{2}+A(N^{*})^{2}+B(I+I^{*})^{2}+(I^{*})^{2}+(R+R^{*})^{2}+(R^{*})^{2}$
for some positive constants $A,B$ to be chosen later. Recalling that $N^{*}=S^{*}+I^{*}+R^{*}$, this is easily seen to be a positive definite quadratic form in $y$ from which we immediately have bounds $c_{1}\left|y\right|^{2}\leq V(y)\leq c_{2}\left|y\right|^{2}$ which shows that $V(y)$ is radially coercive and also reduces our goal of proving $\frac{d}{dt}V(y)\leq-cV(y)$ to proving that $\frac{d}{dt}V(y)\leq-c\left|y\right|^{2}$ for some positive constant $c.$ We compute
(A.5) $\begin{split}\tfrac{d}{dt}V(y)&=-2\gamma(I+I^{*})(\Sigma+I+S^{*}+I^{*})-2\eta I(u+I+S^{*}+I^{*})-2\delta(u+I+S^{*}+I^{*})^{2}\\ &\hskip 28.45274pt+2A\Big{(}(\overline{\mu}-\mu)(u+\tfrac{b}{\delta}+I+R)-(\nu+\delta)\Big{)}(N^{*})^{2}\\ &\hskip 28.45274pt+2B\Big{[}\beta\Big{(}(1-\alpha)(u+\tfrac{b}{\delta})+S^{*}\Big{)}(I+I^{*})^{2}-(\delta+\gamma+\eta)I^{2}-(\delta+\gamma)(I^{*})^{2}\Big{]}\\ &\hskip 28.45274pt+2I^{*}(\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu)IN^{*}-(\nu+\delta)I^{*})\\ &\hskip 28.45274pt+2(R+R^{*})(\gamma(I+I^{*})+\eta I-\delta(R+R^{*}))\\ &\hskip 28.45274pt+2R^{*}(\gamma I^{*}+(\overline{\mu}-\mu)RN^{*}-(\nu+\delta)R^{*})\\ &=\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}L_{1}+L_{2}+L_{3}+L_{4}+L_{5}+L_{6},\end{split}$
where $L_{i}$ is the $i^{\text{th}}$ line of the long, preceding equation. We then bound line-by-line, recalling that the ultimate goal is $\frac{d}{dt}V(y)\leq-c\left|y\right|^{2}$, so we want square terms. We will frequently use the total population bound $\tfrac{b}{\delta}$, and the nonnegativity of each of $I,R,S^{*},I^{*},R^{*}$. In particular, by nonnegativity, any subset of the total population also obeys the total population bound as well. For example, $I+I^{*},\Sigma+\tfrac{b}{\delta}+I+R,N^{*}\leq\tfrac{b}{\delta}.$
With this in mind, from the first line, we drop any cross terms certain to be negative (as well as some of the negative square terms that we won’t need), and then repeatedly use the Cauchy-Schwarz inequality $ab\leq\frac{\varepsilon a^{2}}{2}+\frac{b^{2}}{2\varepsilon}$ for different values of $\varepsilon>0$ to see
(A.6) $\begin{split}L_{1}&\leq-2\delta(\Sigma^{2}+I^{2}+(S^{*})^{2}+(I^{*})^{2})-2(\gamma+\eta+2\delta)\Sigma I-2(\gamma+2\delta)\Sigma I^{*}-4\delta\Sigma S^{*}\\ &\leq-\delta\Sigma^{2}-\Big{(}2\delta-\tfrac{3(\gamma+\nu+2\delta)}{\delta}\Big{)}I^{2}+11\delta(S^{*})^{2}-\Big{(}2\delta-\tfrac{3(\gamma+2\delta)}{\delta}\Big{)}(I^{*})^{2}.\end{split}$
For $L_{2}$, we simply use $\Sigma+\tfrac{b}{\delta}+I+R\leq\tfrac{b}{\delta},$ as well as our assumption that $\frac{b}{\delta}<\frac{\nu+\delta}{\overline{\mu}-\mu},$ to see
(A.7) $\begin{split}L_{2}&\leq 2A\Big{(}(\overline{\mu}-\mu)\tfrac{b}{\delta}-(\nu+\delta)\Big{)}(N^{*})^{2}\leq-2\tilde{A}(N^{*})^{2}\\ &\leq-2\tilde{A}((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2}).\end{split}$
where $\tilde{A}=A((\nu+\delta)-\tfrac{b}{\delta}(\overline{\mu}-\mu))>0$ can be made arbitrarily large by taking $A$ large. Continuing with $L_{3}$, the difficulty is how to handle the triple product term $\Big{(}(1-\alpha)(\Sigma+\tfrac{b}{\delta})+S^{*}\Big{)}(I+I^{*})^{2}$, bearing in mind that we want to keep the full effect of the $(1-\alpha)$ because this appears in $\mathscr{R}_{0}(\tfrac{b}{\delta},0).$ Thus we split up the $(\Sigma+\tfrac{b}{\delta})$ and the $S^{*}$ in the first term and use
(A.8) $\begin{split}(\Sigma+\tfrac{b}{\delta})(I+I^{*})^{2}&\leq(\tfrac{b}{\delta})I^{2}+2(\tfrac{b}{\delta})II^{*}+(\tfrac{b}{\delta})(I^{*})^{2}\\ &\leq\Big{(}(\tfrac{b}{\delta})+\varepsilon\Big{)}I^{2}+C\Big{(}1+\tfrac{1}{\varepsilon}\Big{)}(I^{*})^{2}\end{split}$
where $\varepsilon>0$ will be specified later, and $C$ is some positive constant comprised of the other parameters (in what follows, this constant may change from line-to-line but it will always be a finite, positive constant). Next, we use
(A.9) $\begin{split}S^{*}(I+I^{*})^{2}&\leq(\tfrac{b}{\delta})S^{*}(I+I^{*})\\ &\leq\varepsilon I^{2}+C(1+\tfrac{1}{\varepsilon})(S^{*})^{2}+C(I^{*})^{2}.\end{split}$
Combining (A.8) and (A.9) with the other terms in line 3 of (A.5), we have
(A.10) $L_{3}\leq 2B\Big{[}(1-\alpha)\Big{(}(\tfrac{b}{\delta})+2\varepsilon)-(\gamma+\delta+\eta)\Big{]}I^{2}+C(1+\tfrac{1}{\varepsilon})((S^{*})^{2}+(I^{*})^{2}).$
For line 4 in (A.5), we use $I,I+I^{*}\leq\tfrac{b}{\delta}$ so that all terms are simply quadratic in $S^{*},I^{*},R^{*}$ whereupon
(A.11) $L_{4}\leq C((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2}).$
For line 5 in (A.5), the only bothersome terms are those involving $I$ and $R$, and we have
(A.12) $L_{5}\leq\tfrac{2(\gamma+\eta)}{\delta}I^{2}-\delta R^{2}+C((I^{*})^{2}+(R^{*})^{2}).$
And finally, for line 6 of (A.5), once we bound $R\leq\tfrac{b}{\delta}$, all terms are quadratic in $S^{*},I^{*},R^{*}$ again so
(A.13) $L_{6}\leq C((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2}).$
Finally, combining the six bounds (A.6), (A.7),(A.10),(A.11),(A.12),(A.13), we arrive at
(A.14) $\begin{split}\tfrac{d}{dt}V(y)&\leq-\Big{(}2\delta-2B\Big{[}(1-\alpha)\Big{(}(\tfrac{b}{\delta})+2\varepsilon)-(\gamma+\delta+\eta)\Big{]}-\tfrac{3(\gamma+\nu+2\delta)+2(\gamma+\eta)}{\delta}\Big{)}I^{2}\\ &\hskip 28.45274pt-\delta(\Sigma^{2}+R^{2})+\Big{[}C\Big{(}1+\tfrac{1}{\varepsilon}\Big{)}-2\tilde{A}\Big{]}((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2})\\ &=\Big{(}2\delta+\tfrac{2B}{\gamma+\eta+\delta}\Big{[}1-\mathscr{R}_{0}(\tfrac{b}{\delta},0)-\tfrac{2\varepsilon}{\gamma+\eta+\delta}\Big{]}-\tfrac{3(\gamma+\nu+2\delta)+2(\gamma+\eta)}{\delta}\Big{)}I^{2}\\ &\hskip 28.45274pt-\delta(\Sigma^{2}+R^{2})+\Big{[}C\Big{(}1+\tfrac{1}{\varepsilon}\Big{)}-2\tilde{A}\Big{]}((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2}).\end{split}$
Since we have assumed $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$, we can take $\varepsilon>0$ small enough that $1-\mathscr{R}_{0}(\tfrac{b}{\delta},0)-\tfrac{2\varepsilon}{\gamma+\eta+\delta}>0$, and then we can choose $B>0$ so that
$\tfrac{2B}{\gamma+\eta+\delta}\Big{[}1-\mathscr{R}_{0}(\tfrac{b}{\delta},0)-\tfrac{2\varepsilon}{\gamma+\eta+\delta}\Big{]}-\tfrac{3(\gamma+\nu+2\delta)+2(\gamma+\eta)}{\delta}=0.$
Finally, having fixed $\varepsilon>0$, we simply take $A$ large enough so that $C(1+\tfrac{1}{\varepsilon})-2\tilde{A}<0$, and we arrive at $\tfrac{d}{dt}V(y)\leq-c\left|y\right|^{2}$ for a positive constant $c$, whereupon the DFE $(s,s^{*})=(\tfrac{b}{\delta},0)$ is globally asymptotically stable.
As stated before, we refer the reader to [39] for other cases. ∎
Chloe Ngo, Honors College, University of Oklahoma, Norman, OK, USA; e-mail: chloe.ngo@ou.edu
Christian Parkinson, Department of Mathematics & Department of Computational Mathematics, Science and Engineering, Michigan State University, East Lansing, MI, USA; e-mail: chparkin@msu.edu
Weinan Wang, Department of Mathematics, University of Oklahoma, Norman, OK, USA; e-mail: ww@ou.edu

### v2

@@
 We derive the reproductive ratio for our model and provide stability analysis for the disease-free equilibria.
-We then propose a control scenario wherein a policy-maker with access to control variables representing disease prevention mandates, treatment efforts, and educational campaigns aimed at encouraging compliance minimizes a cost functional incorporating several cost concerns.
-We characterize optimal controls via the Pontryagin optimality principle and present simulations which demonstrate the behavior of the control maps in several different parameter regimes.
+We then propose an optimal control scenario wherein a policy-maker with access to control variables representing disease prevention mandates, treatment efforts, and educational campaigns aimed at encouraging compliance minimizes a cost functional incorporating several cost concerns.
+Via careful analysis of the control-to-state map, we are able to prove existence of optimal controls.
+Our proof applies to dynamics which can be nonlinear in the control variables and general cost functionals including the case of $L^{1}$ control costs.
+We numerically resolve optimal strategies using the sequential quadratic Hamiltonian method, a relatively new numerical method for optimal control which is easy to implement and has good convergence theory, as we demonstrate.
+We test our model in several parameter regimes with specific interest in observing how the policy-maker’s optimal strategies depend on their particular preferences which are expressed via design of different cost functionals.
 1. Introduction
@@
 Incorporation and assessment of the effects of heterogeneous behavior into epidemiological models is becoming increasingly popular (see, for example, [5, 26, 32, 39, 41], or [50] where the authors propose a rather general framework similar to the modeling we use below, though with linear transfer between behavior classes).
-Likewise, there are several recent studies of optimal control for epidemiological models in different modeling scenarios, such as optimal treatment or vaccination [6, 13, 26, 50], optimal quarantine under a variety of limitations [3, 15], and assessment of the effects of contact heterogeneity [14] or the use of different types of cost functionals [44].
+Likewise, there are several recent studies of optimal control for epidemiological models in different modeling scenarios, such as optimal treatment or vaccination [1, 7, 23, 31, 59], optimal quarantine under a variety of limitations [4, 17], and assessment of the effects of contact heterogeneity [16] or the use of different types of cost functionals [51].
 However, very few studies include both optimal control for epidemiology and mechanistic modeling of human behavior.
@@
 Other studies incorporate control and human behavior more implicitly, by, for example, allowing infection rates to depend on ambient dynamics under the assumption that people will change their behavior based on local disease effects [42] or media coverage [20].
-In this manuscript, we develop a compartmental epidemiological model which incorporates human behavior in an explicit, mechanistic manner.
+1.1. Our Contribution.
+We develop a compartmental epidemiological model which incorporates human behavior in an explicit, mechanistic manner.
 We then consider optimal control of the model and observe the differences in optimal control strategies in different parameter regimes.
+In particular, we prove existence of optimal controls in a generic scenario wherein (i) controls appear in the model nonlinearly, (ii) control costs are not required to be smooth or strongly convex (e.g.
+$L^{1}$ control costs are allowable), and (iii) the admissible control set contains discontinuous control maps.
+With small tweaks, our existence proof will go through for many biological models where (i)-(iii) are physically preferable, but are commonly relaxed (e.g.
+by the restriction to control affine systems and use of $L^{2}$ control costs) for ease of analysis.
+We resolve optimal controls via the sequential quadratic Hamiltonian (SQH) method [10].
+This is relatively new algorithm for optimal control which is black-box free and has a reasonable convergence theory, as demonstrated below.
+This distinguishes SQH from open source large scale optimization packages (used, for example, in [7, 17]) which are explainable to experts, but often operate as black-boxes for the end user, and also from the popular forward-backward sweep method (used, for example, in [1, 20, 23]) which is easy to implement but does not have a robust convergence theory.
+To the authors’ knowledge, this method has not yet appeared in the epidemiology literature, though it does appear in the preprint [43] in the context of control for a partial differential equation model for epidemics.
+Thus, the contribution of this manuscript goes further than the particular model we analyze, but also includes a general existence framework for optimal controls that applies to biologically realistic models, and the introduction to the epidemiology literature of a numerical method for optimal control which may be preferable to those commonly used.
 The remainder of the manuscript is organized as follows.
-In section 2, we introduce our model, derive the reproductive ratio, and discuss stability of disease-free equilibria (though a full proof of our stability theorem is relegated to appendix A).
-In section 3, we introduce the optimal control problem that we will solve and give a characterization of the optimal control maps via the Pontryagin optimality principle.
-In section 4, we discuss the results of simulation of our model in different parameter regimes.
+In section 2, we introduce our model, derive the reproductive ratio, and discuss stability of disease-free equilibria.
+In section 3, we introduce the optimal control problem that we will solve and prove existence of optimal controls for a reasonably generic cost functional.
+In particular, our existence proof allows for $L^{1}$ control cost and discontinuous controls, which are physically meaningful for our model.
+In section 4, we describe the application of the SQH method [10] to our problem, and prove convergence of the algorithm.
+In section 5, we discuss the results of simulations of our model in different parameter regimes.
+Specifically, we make some qualitative observations regarding the optimal control strategies that result from different policy-maker preferences.
 Finally, section 6 includes a brief conclusion and discussion of avenues for future work.
@@
 With this assumption, we propose the following model:
-(2.2) $\begin{split}\frac{dS}{dt}&=(1-\xi)b-\beta(1-\alpha(t))S(I+I^{*})-(\overline{\mu}-\mu(t))SN^{*}+\nu(t)S^{*}-\delta S,\\ \frac{dI}{dt}&=\beta(1-\alpha(t))S(I+I^{*})-(\gamma+\eta(t))I-(\overline{\mu}-\mu(t))IN^{*}+\nu(t)I^{*}-\delta I,\\ \frac{dR}{dt}&=(\gamma+\eta(t))I-(\overline{\mu}-\mu(t))RN^{*}+\nu(t)R^{*}-\delta R,\\ \frac{dS^{*}}{dt}&=\xi b-\beta S^{*}(I+I^{*})+(\overline{\mu}-\mu(t))SN^{*}-\nu(t)S^{*}-\delta S^{*},\\ \frac{dI^{*}}{dt}&=\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu(t))IN^{*}-\nu(t)I^{*}-\delta I^{*},\\ \frac{dR^{*}}{dt}&=\gamma I^{*}+(\overline{\mu}-\mu(t))RN^{*}-\nu(t)R^{*}-\delta R^{*}.\end{split}$
+(2.2) $\begin{split}\frac{dS}{dt}&=(1-\xi)b-\beta(1-\alpha(t))S((1-\alpha(t))I+I^{*})-(\mu^{*}-\mu(t))SN^{*}+\nu(t)S^{*}-\delta S,\\ \frac{dI}{dt}&=\beta(1-\alpha(t))S((1-\alpha(t))I+I^{*})-(\gamma+\eta(t))I-(\mu^{*}-\mu(t))IN^{*}+\nu(t)I^{*}-\delta I,\\ \frac{dR}{dt}&=(\gamma+\eta(t))I-(\mu^{*}-\mu(t))RN^{*}+\nu(t)R^{*}-\delta R,\\ \frac{dS^{*}}{dt}&=\xi b-\beta S^{*}((1-\alpha(t))I+I^{*})+(\mu^{*}-\mu(t))SN^{*}-\nu(t)S^{*}-\delta S^{*},\\ \frac{dI^{*}}{dt}&=\beta S^{*}((1-\alpha(t))I+I^{*})-\gamma I^{*}+(\mu^{*}-\mu(t))IN^{*}-\nu(t)I^{*}-\delta I^{*},\\ \frac{dR^{*}}{dt}&=\gamma I^{*}+(\mu^{*}-\mu(t))RN^{*}-\nu(t)R^{*}-\delta R^{*}.\end{split}$
 Here the birth rate $b$, death rate $\delta$, infectivity rate $\beta$ and recovery rate $\gamma$ have the same interpretation as in (2.1).
 New parameters are as follows: $\xi\in[0,1]$ is the probability with which newly introduced members of the population are noncompliant, and hence the birth rate $b$ is split between the $S$ and $S^{*}$ equations with contributions $(1-\xi)b$ and $\xi b$ respectively.
-The total noncompliant population is defined by $N^{*}=S^{*}+I^{*}+R^{*}$ and is involved in new mass action terms which cause transfer from $S,I,R$ to $S^{*},I^{*},R^{*}$ with baseline infectivity rate $\overline{\mu}$ (the assumption being that anytime a compliant individual comes in contact with a noncompliant individual, they have some chance of becoming noncompliant).
+The total noncompliant population is defined by $N^{*}=S^{*}+I^{*}+R^{*}$ and is involved in new mass action terms which cause transfer from $S,I,R$ to $S^{*},I^{*},R^{*}$ with baseline infectivity rate $\mu^{*}$ (the assumption being that anytime a compliant individual comes in contact with a noncompliant individual, they have some chance of becoming noncompliant).
 The remaining parameters $\alpha,\eta,\mu,\nu$ represent the policy-maker’s attempts to stop the spread of the disease and/or noncompliance.
@@
   • (1)
-    $\alpha(\cdot)\in[0,1]$ - reduction in infectivity among interactions with the compliant susceptible population.
+    $\alpha(\cdot)\in[0,\overline{\alpha}]$ - reduction in infectivity or mixing for the compliant population.
+Here $\overline{\alpha}\in[0,1]$ is the maximum decrease in mixing / infectivity that can be achieved through the use of NPIs.
+Note, $\alpha(t)$ appears in the mass action terms representing spread of infections; in these terms, the compliant subpopulations $S,I$ are multiplied by $(1-\alpha(t))$, while the noncompliant subpopulations $S^{*},I^{*}$ are not.
   • (2)
@@
     $\mu(\cdot)\in[0,\overline{\mu}]$ - reduction in spread of noncompliance due to public health information initiatives.
+Here $\overline{\mu}\in[0,\mu^{*}]$ is the maximum achievable decrease in the spread rate for noncompliance.
   • (4)
@@
 Here $\overline{\nu}$ is the maximum achievable recovery rate from noncompliance.
+We define $U_{\text{ad}}=[0,\overline{\alpha}]\times[0,\overline{\eta}]\times[0,\mu^{*}]\times[0,\overline{\nu}]$ to be our set of admissible control actions.
 The uncontrolled scenario would be $\alpha(\cdot),\eta(\cdot),\mu(\cdot),\nu(\cdot)\equiv 0$, and the maximally controlled scenario would have each control variable identically equal to its maximum value.
@@
 Before that, we cover some standard considerations regarding compartmental epidemiological models; namely, nonnegativity of solutions, derivation of the reproductive ratio, and stability of disease-free equilibria.
-For the remainder of this section, we treat $\alpha,\eta,\mu,\nu$ as fixed, nonnegative, constant values.
-We will go through this quickly; much of the same work is included in more detail in [39], albeit for a slightly different system of equations.
-Because we will use the next generation matrix method [19, 46, 47] to derive the reproductive ratio for our system, we write (2.1) abstractly as
-(2.3) $x^{\prime}=\mathcal{F}(x)+\mathcal{V}^{+}(x)-\mathcal{V}^{-}(x)$
-where we reorder the variables $x=(I,I^{*},S,R,S^{*},R^{*})$ by bringing the infected compartments to the front, and define
-(2.4) $\mathcal{F}(x)=\begin{bmatrix}\beta(1-\alpha)S(I+I^{*})\\ \beta S^{*}(I+I^{*})\\ 0\\ 0\\ 0\\ 0\end{bmatrix}$
-and
-(2.5) $\mathcal{V}^{+}(x)=\begin{bmatrix}\nu I^{*}\\ (\overline{\mu}-\mu)IN^{*}\\ (1-\xi)b+\nu S^{*}\\ (\gamma+\eta)I+\nu R^{*}\\ \xi b+(\overline{\mu}-\mu)SN^{*}\\ \gamma I^{*}+(\overline{\mu}-\mu)RN^{*}\end{bmatrix},\,\,\,\,\,\,\,\,\,\,\mathcal{V}^{-}(x)=\begin{bmatrix}(\gamma+\eta+\delta+(\overline{\mu}-\mu)N^{*})I\\ (\gamma+\nu+\delta)I^{*}\\ (\beta(1-\alpha)(I+I^{*})+(\overline{\mu}-\mu)N^{*}+\delta)S\\ ((\overline{\mu}-\mu)N^{*}+\delta)R\\ (\beta(I+I^{*})+\nu+\delta)S^{*}\\ (\nu+\delta)R^{*}\end{bmatrix}$.
-Here $\mathcal{F}$ includes all terms which correspond to adding new infections into the system, $\mathcal{V}^{+}$ includes all terms corresponding to in-flow to a compartment caused by anything other than new infections, and $\mathcal{V}^{-}$ includes all terms corresponding to outflow from a compartment for any reason.
-We note in particular that $\mathcal{F},\mathcal{V}^{+},\mathcal{V}^{-}\geq 0$ when $S,I,R,S^{*},I^{*},R^{*}\geq 0$.
+We note that this work is already contained in [45], except in that manuscript $\eta=0$.
+Because $\eta>0$ causes an asymmetry in the way $I$ and $I^{*}$ are treated, the derivations here are a bit messier, though no more difficult.
+Thus to avoid duplication, we will cover them tersely and refer the reader to [45] for proofs and expanded reasoning.
 Proposition 2.1.
-Assuming nonnegative initial conditions, (2.2) admits a unique, global-in-time, nonnegative solution.
-Proof.
-A unique local solution to (2.2) exists as consequence of the Picard-Lindelöf theorem since the nonlinearities are Lipschitz.
-Nonnegativity is preserved as long as the solution exists because, looking at each component $x_{j}$ of (2.3), we have $x_{j}^{\prime}\geq 0$ when $x_{j}=0$ since the $i^{\text{th}}$ component of $\mathcal{V}^{-}$ is proportional to $x_{i}$.
-Given nonnegativity, we see that solutions cannot blow up because the total population $N_{\text{total}}=S+I+R+S^{*}+I^{*}+R^{*}$ satisfies
+Assuming nonnegative initial conditions and $(\alpha,\eta,\mu,\nu)\in L^{\infty}([0,T];U_{\text{ad}})$, (2.2) admits a unique, nonnegative, a.e. solution on the time interval $[0,T].$
+The existence and uniqueness follow directly from Caratheodory’s theorem [49, Ch.
+1.6].
+The preservation of nonnegativity in proposition 2.1 follows since $\frac{dX}{dt}|_{X=0}\geq 0$ whenever all components are nonnegative, for each $X\in\{S,I,R,S^{*},I^{*},R^{*}\}.$ We also note that there is a total population bound.
+Indeed, defining $N_{\text{total}}=S+I+R+S^{*}+I^{*}+R^{*}$, we see
 (2.3) $N_{\text{total}}^{\prime}=b-\delta N_{\text{total}}\,\,\,\,\,\implies\,\,\,\,\,N_{\text{total}}(t)=\frac{b}{\delta}+\bigg(N_{\text{total}}(0)-\frac{b}{\delta}\bigg)e^{-\delta t}\leq\max\left\{N_{\text{total}}(0),\frac{b}{\delta}\right\}$
-which serves as an upper bound on each of the subpopulations. ∎
-We can assume without loss of generality that $N_{\text{total}}(0)=1$, whereupon alternately setting $\tfrac{b}{\delta}\geq 1$ or $\tfrac{b}{\delta}<1$ would correspond to situations where the population is increasing or where the population is strictly decreasing, respectively.
-For ease of analysis, we will always assume that $\tfrac{b}{\delta}\geq 1=N(0)$, so that $\tfrac{b}{\delta}$ serves as a bound on the total population.
-A disease-free equilibrium (DFE) point for (2.2) is an equilibrium point wherein $I=I^{*}=0$.
-By adding the equations for $R,R^{*}$ together, we see that any DFE will also have $R=R^{*}=0$, whereupon $x=(0,0,s,0,s^{*},0)$ is a DFE if and only if
-(2.7) $\begin{split}(1-\xi)b-(\overline{\mu}-\mu)ss^{*}+\nu s^{*}-\delta s&=0,\\ \xi b+(\overline{\mu}-\mu)ss^{*}-\nu s^{*}-\delta s^{*}&=0.\end{split}$
-Because we are looking for physically meaningful solutions, we only consider values $0\leq s,s^{*}\leq\frac{b}{\delta}$.
-With this restriction in mind, we specify the different parameter regimes which lead to existence of different DFEs.
-If we assume $\xi=0$ so that all newly introduced members of the population are compliant, there are two solutions which can be physically meaningful:
-(2.8) $\begin{split}s_{1}&=\tfrac{b}{\delta},\\ s^{*}_{1}&=0,\end{split}\hskip 56.9055pt\text{ and }\hskip 56.9055pt\begin{split}s_{2}&=\tfrac{\nu+\delta}{\overline{\mu}-\mu},\\ s^{*}_{2}&=\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu},\end{split}$
+We can assume without loss of generality that $N_{\text{total}}(0)=1$, whereupon alternately setting $\tfrac{b}{\delta}\geq 1$ or $\tfrac{b}{\delta}<1$ would correspond to situations where the population is increasing or where the population is strictly decreasing.
+For the remainder of this section, we treat $\alpha,\eta,\mu,\nu$ as fixed, nonnegative, constant values for the sake of finding disease free equilibrium (DFE) points.
+The DFEs for out system are equilibrium points $x^{*}=(s,i,r,s^{*},i^{*},r^{*})$ with $i=i^{*}=r=r^{*}=0$ and $s,s^{*}\geq 0$.
+There are two DFEs if $\xi=0$ (meaning that all newly introduced members of the population are compliant).
+These are given by
+(2.4) $\begin{split}s_{1}&=\tfrac{b}{\delta},\\ s^{*}_{1}&=0,\end{split}\hskip 56.9055pt\text{ and }\hskip 56.9055pt\begin{split}s_{2}&=\tfrac{\nu+\delta}{\mu^{*}-\mu},\\ s^{*}_{2}&=\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu},\end{split}$
 The first is always physically meaningful and represents the “best case scenario” for the policy-maker wherein the whole population is compliant.
-The second is physically meaningful and distinct only if $\tfrac{b}{\delta}>\frac{\nu+\delta}{\overline{\mu}-\mu}.$ Since $\mu\in[0,\overline{\mu}]$, the denominator of $\frac{\nu+\delta}{\overline{\mu}-\mu}$ is nonnegative, so in the case that $\mu\to\overline{\mu}$, the second DFE is no longer meaningful.
-If instead $\xi\in(0,1]$ so that some portion of the newly introduced population is noncompliant, (2.7) has a unique physically meaningful solution
-(2.9) $\begin{split}s_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}+\tfrac{\delta+\nu}{\overline{\mu}-\mu}-\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\overline{\mu}-\mu}\right)^{2}+\tfrac{4\xi b}{\overline{\mu}-\mu}}\right),\\ s^{*}_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\overline{\mu}-\mu}+\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\overline{\mu}-\mu}\right)^{2}+\tfrac{4\xi b}{\overline{\mu}-\mu}}\right).\end{split}$
-In the limit as $\xi\to 0$, we find $(s_{3},s_{3}^{*})\to(s_{1},s_{1}^{*})$ when $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\overline{\mu}-\mu}$, and $(s_{3},s_{3}^{*})\to(s_{2},s_{2}^{*})$ when $\tfrac{b}{\delta}>\frac{\nu+\delta}{\overline{\mu}-\mu}$.
+The second is physically meaningful and distinct only if $\tfrac{b}{\delta}>\frac{\nu+\delta}{\mu^{*}-\mu}.$ Since $\mu\in[0,\mu^{*}]$, the denominator of $\frac{\nu+\delta}{\mu^{*}-\mu}$ is nonnegative, so in the case that $\mu\to\mu^{*}$, the second DFE is no longer meaningful.
+If instead $\xi\in(0,1]$ so that some portion of the newly introduced population is noncompliant, there is a unique DFE
+(2.5) $\begin{split}s_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}+\tfrac{\delta+\nu}{\mu^{*}-\mu}-\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\mu^{*}-\mu}\right)^{2}+\tfrac{4\xi b}{\mu^{*}-\mu}}\right),\\ s^{*}_{3}&=\tfrac{1}{2}\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\mu^{*}-\mu}+\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\delta+\nu}{\mu^{*}-\mu}\right)^{2}+\tfrac{4\xi b}{\mu^{*}-\mu}}\right).\end{split}$
+In the limit as $\xi\to 0$, we find $(s_{3},s_{3}^{*})\to(s_{1},s_{1}^{*})$ when $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\mu^{*}-\mu}$, and $(s_{3},s_{3}^{*})\to(s_{2},s_{2}^{*})$ when $\tfrac{b}{\delta}>\frac{\nu+\delta}{\mu^{*}-\mu}$.
 A special case of (2.5) is the “worst case scenario” for the policy-maker wherein $\xi=1$ and $\nu=0$, meaning that all newly introduced members of the population are noncompliant and noncompliance is a permanent state.
 In this case, the DFE is $s_{3}=0,s^{*}_{3}=\tfrac{b}{\delta}$.
-Given this, we can compute the reproductive ratio $\mathscr{R}_{0}(s,s^{*})$ corresponding to the DFE $x_{s,s^{*}}=(s,s^{*})$.
-Roughly speaking, $\mathscr{R}_{0}(s,s^{*})$ can be interpreted as the expected number of secondary infections that result from any one infection.
-This can be computed via the next generation matrix method [19, 46, 47]: it is the spectral radius of the matrix $FV^{-1}$ where $F$ and $V$ are, respectively, the leading principal submatrices of $D\mathcal{F}(x_{s,s^{*}})$ and $D\mathcal{V}^{-}(x_{s,s^{*}})-D\mathcal{V}^{+}(x_{s,s^{*}})$ corresponding to the infected compartments, where $\mathcal{F},\mathcal{V}^{+},\mathcal{V}^{-}$ are as in (2.4),(2.5).
-In our case, these are
-(2.10) $F=\begin{bmatrix}\beta(1-\alpha)s&\beta(1-\alpha)s\\ \beta s^{*}&\beta s^{*}\end{bmatrix},\hskip 56.9055ptV=\begin{bmatrix}\gamma+\eta+\delta+(\overline{\mu}-\mu)s^{*}&-\nu\\ -(\overline{\mu}-\mu)s^{*}&\gamma+\nu+\delta\end{bmatrix}$.
-Because $F$ is rank 1, the spectral radius $\rho(FV^{-1})=\mathscr{R}_{0}(s,s^{*})$ can easily be computed by hand.
-We write it in two different equivalent ways to point out behavior with respect to changes in different parameters:
-(2.11) $\begin{split}\mathscr{R}_{0}(s,s^{*})&=\frac{\beta\Big{[}(1-\alpha)\Big{(}\gamma+\nu+\delta+(\overline{\mu}-\mu)s^{*}\Big{)}s+\Big{(}\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*}\Big{)}s^{*}\Big{]}}{(\gamma+\delta)(\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*})+\eta\nu}\\ &=\frac{\beta\Big{[}(1-\alpha)\Big{(}1-\tfrac{\eta}{\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*}}\Big{)}s+s^{*}\Big{]}}{\gamma+\delta+\frac{\eta\nu}{\gamma+\eta+\nu+\delta+(\overline{\mu}-\mu)s^{*}}}.\end{split}$
-The second representation lends itself to a reasonably natural interpretation, with the main complicating factor being the difference in recovery rates between the compliant and noncompliant populations due to the compliant population seeking treatment represented by $\eta$.
+The reproductive ratio $\mathscr{R}_{0}(s,s^{*})$ corresponding to a DFE $x_{s,s^{*}}=(s,0,0,s^{*},0,0)$ can be computed by the next generation matrix method [21, 54, 55]:
+(2.6) $\mathscr{R}_{0}(s,s^{*})=\frac{\beta\Big[(1-\alpha)^{2}\Big(1+\tfrac{\alpha\mu s^{*}/(1-\alpha)-\eta}{\gamma+\eta+\nu+\delta+(\mu^{*}-\mu)s^{*}}\Big)s+\Big(1-\tfrac{\alpha\nu}{\gamma+\eta+\nu+\delta+(\mu^{*}-\mu)s^{*}}\Big)s^{*}\Big]}{\gamma+\delta+\frac{\eta\nu}{\gamma+\eta+\nu+\delta+(\mu^{*}-\mu)s^{*}}}.$
+This has a reasonably natural interpretation, with the main complicating factor being the difference in recovery rates between the compliant and noncompliant populations due to the compliant population seeking treatment represented by $\eta$.
 In the worst case scenario, where $\xi=1,\nu=0$ so that the DFE is $(s,s^{*})=(0,\tfrac{b}{\delta})$, we find $\mathscr{R}_{0}(0,\tfrac{b}{\delta})=\frac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}.$ This is the reproductive ratio for the basic SIR model (2.1), which signifies that in this worst case scenario, the policy-maker’s control efforts do nothing to stop the spread of the disease.
-In the best case scenario, $\xi=0$ and $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\overline{\mu}-\mu}$ so that the only DFE is $(s,s^{*})=(\tfrac{b}{\delta},0)$ and we find that $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\frac{\beta(1-\alpha)}{\gamma+\eta+\delta}.$ This decreases the reproductive ratio when compared with the basic SIR model in two ways: $(1-\alpha)$ in the numerator results from the reduction in infectivity among the compliant susceptible population and the $\eta$ in the denominator results from the increased recovery rate among the compliant infected population.
-In general, $\mathscr{R}_{0}(s,s^{*})$ can then be viewed as a very complicated interpolation between these two endpoints which is illuminated a bit in the second line of (2.11) where the terms in front of $s$ in the numerator can be seen to represent reduction in disease spread among the compliant population, and that $s^{*}$ has no such modifications represents inefficacy of control efforts among the noncompliant population.
+In the best case scenario, $\xi=0$ and $\tfrac{b}{\delta}\leq\frac{\nu+\delta}{\mu^{*}-\mu}$ so that the only DFE is $(s,s^{*})=(\tfrac{b}{\delta},0)$ and we find that $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\frac{\beta(1-\alpha)^{2}}{\gamma+\eta+\delta}.$ This decreases the reproductive ratio when compared with the basic SIR model in two ways: $(1-\alpha)^{2}$ in the numerator results from the reduction in infectivity among the compliant populations and the $\eta$ in the denominator results from the increased recovery rate among the compliant infected population.
+In general, $\mathscr{R}_{0}(s,s^{*})$ can then be viewed as a very complicated interpolation between the two points wherein the compliant susceptible population $s$ is receiving some reduction in infectivity signified by the $(1-\alpha)^{2}$, whereas the noncompliant susceptible population $s^{*}$ receives no such benefit.
 We make some natural observations regarding $\mathscr{R}_{0}(s,s^{*})$.
@@
 The two points can be interpreted as follows: (i) more control leads to a lower rate of disease spread, and (ii) when we are at the equilibrium population level of $\tfrac{b}{\delta}$, more noncompliance leads to a higher rate of disease spread.
-Proof.
-That $\mathscr{R}_{0}(s,s^{*})$ is decreasing in $\alpha$ when other values are held constant is obvious directly from (2.11) since $\mathscr{R}_{0}(s,s^{*})$ is linear in $\alpha$.
-That $\mathscr{R}_{0}(s,s^{*})$ is decreasing in $\mu$ can be seen from the second line (2.11) since increasing $\mu$ increases each of the nested fractions, each of which has the effect of decreasing $\mathscr{R}_{0}(s,s^{*}).$ The remaining statements in (i), (ii) are not so obvious directly from the formula, but follow from the observation that $z\mapsto\frac{c_{1}+c_{2}z}{c_{3}+c_{4}z}$ is increasing for $z\geq 0$ when $c_{2}c_{3}\geq c_{1}c_{4}$ and decreasing for $z\geq 0$ when $c_{2}c_{3}\leq c_{1}c_{4}$.
-For (i), alternately writing $\mathscr{R}(s,s^{*})$ in this form where the role of $z$ is played by $\eta$ or $\nu$, the inequalities proving that $\mathscr{R}_{0}(s,s^{*})$ is decreasing boils down to
-$(\gamma+\delta)(\gamma+\nu+\delta+(\overline{\mu}-\mu)s^{*})s^{*}$ $\leq(\gamma+\delta+\nu)(\gamma+\nu+\delta+(\overline{\mu}-\mu)s^{*})((1-\alpha)s+s^{*}),$
-$(\gamma+\delta)(\gamma+\eta+\delta+(\overline{\mu}-\mu)s^{*})$ $\leq(\gamma+\eta+\delta)\Big{(}\gamma+\delta+(\overline{\mu}-\mu)s^{*}+\tfrac{\eta s^{*}}{(1-\alpha)s+s^{*}}\Big{)},$
-respectively. Each of these is verified simply by checking that the right-hand side has all the same terms as the left hand side plus a few extra positive terms.
-For part (ii), if $s^{*}\in[0,\tfrac{b}{\delta}]$, we see
-(2.12) $\mathscr{R}_{0}(\tfrac{b}{\delta}-s^{*},s^{*})=\frac{\beta\Big{[}\tfrac{b}{\delta}(\gamma+\nu+\delta)+\Big{(}\tfrac{b}{\delta}(\overline{\mu}-\mu)+\eta\Big{)}s^{*}\Big{]}}{\Big{(}(\gamma+\delta)(\gamma+\eta+\nu+\delta)+\eta\nu\Big{)}+(\gamma+\delta)(\overline{\mu}-\mu)s^{*}}.$
-That this is increasing in $s^{*}$ follows from the same reasoning with the inequality flipped. ∎
-To conclude this section, we state stability results for our DFE. We state the theorem here, but relegate the proof to an appendix since it requires a long computation.
+We omit the proof because it is elementary, if tedious.
+It follows most readily from the observation that, the function $z\mapsto\frac{c_{1}+c_{2}z}{c_{3}+c_{4}z}$ in increasing for $z\geq 0$ when $c_{2}c_{3}\geq c_{1}c_{4}$, and decreasing for $z\geq 0$ when $c_{2}c_{3}\leq c_{1}c_{4}$.
+Clearing the nested denominators, one can write $\mathscr{R}_{0}(s,s^{*})$ in this form with $\alpha,\eta,\mu,\nu$ alternately playing the role of $z$, and verify the correct inequality.
+To conclude this section, we state stability results for our DFE.
 Theorem 2.3.
@@
   • (i)
-    If $\xi=0$ and $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$, then $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$ and unstable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)>1$.
+    If $\xi=0$ and $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\mu^{*}-\mu}$, then $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$ and unstable when $\mathscr{R}_{0}(\tfrac{b}{\delta},0)>1$.
   • (ii)
-    If $\xi=0$ and $\tfrac{b}{\delta}>\tfrac{\nu+\delta}{\overline{\mu}-\mu}$, then $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})<1$ and unstable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})>1$
+    If $\xi=0$ and $\tfrac{b}{\delta}>\tfrac{\nu+\delta}{\mu^{*}-\mu}$, then $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})$ is asymptotically stable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})<1$ and unstable when $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})>1$
   • (iii)
@@
 Remark.
-In each case, that the condition on the reproductive ratio determines local asymptotic stability or instability follows directly from [47, Theorem 2] once the assumptions of that theorem are verified.
+In each case, that the condition on the reproductive ratio determines local asymptotic stability or instability follows directly from [55, Theorem 2] once the assumptions of that theorem are verified.
 However, the stability is actually global, in the sense that under the conditions of each statement, any solution of (2.2) will tend toward the given equilibrium point regardless of whether the initial condition is nearby the equilibrium point.
-This global version can be proven using appropriately constructed Lyapunov functions as demonstrated in the appendix and in [39].
+This global version can be proven using appropriately constructed Lyapunov functions.
+We omit the proof because it is quite long and technical, and is contained in [45].
 3. Optimal Control of our Model
 […]
-With this notation, we define the marginal cost function,
-(3.1) $r(x,u)=c_{1}(I+I^{*})+c_{2}N^{*}+\frac{c_{3}}{2}\alpha^{2}+\frac{c_{4}}{2}\eta^{2}+\frac{c_{5}}{2}\mu^{2}+\frac{c_{6}}{2}\nu^{2},$
-where, as above, $N^{*}=S^{*}+I^{*}+R^{*}$. Our cost functional is then
-(3.2) $J(x,u)=\int^{T}_{0}r(x(t),u(t))dt.$
-In (3.1), $c_{1},\ldots,c_{6}$ are positive weights indicating the preferences of the policy-maker.
-For example, a more public-health oriented policy-maker would take $c_{1}$ larger, whereas an economically stingy policy-maker would increase $c_{3},c_{4},c_{5},c_{6}$.
+With this notation, we define a generic cost functional
+(3.1) $J(x,u)=\int^{T}_{0}(r(x(t))+\ell(u(t)))dt$
+where $r:{\mathbb{R}}^{6}_{\geq 0}\to{\mathbb{R}}$ and $\ell:{\mathbb{R}}^{4}\to{\mathbb{R}}$.
+We state assumptions on $r,\ell$ in theorem 3.1, but we make special note that we do not require strict or strong convexity.
+The exact choice of $r,\ell$ indicate the preferences of the policy-maker.
+For example, a more public-health oriented policy-maker would choose $r$ which assigns large values to states with higher $I$ and $I^{*}$, whereas an economically stingy policy-maker would choose $\ell$ to take large values for larger values of the controls.
 From the above, our admissible control set is
-(3.3) $U_{\text{ad}}=\{(\alpha,\eta,\mu,\nu)\in{\mathbb{R}}^{4}\,:\,0\leq\alpha\leq 1,\,0\leq\eta\leq\overline{\eta},\,0\leq\mu\leq\overline{\mu},\,0\leq\nu\leq\overline{\nu}\}$
+(3.2) $U_{\text{ad}}=\{(\alpha,\eta,\mu,\nu)\in{\mathbb{R}}^{4}\,:\,0\leq\alpha\leq\overline{\alpha},\,0\leq\eta\leq\overline{\eta},\,0\leq\mu\leq\overline{\mu},\,0\leq\nu\leq\overline{\nu}\}$
 where $\overline{\eta},\overline{\mu},\overline{\nu}$ are fixed positive constants. We then consider control functions in
-(3.4) $\mathscr{U}_{\text{ad}}=\{u(\cdot)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(\alpha(\cdot),\eta(\cdot),\mu(\cdot),\nu(\cdot)):[0,T]\to U_{\text{ad}}\,\,|\,\,u\text{ measurable}\}.$
-Our optimal control problem is then
+(3.3) $\mathscr{U}_{\text{ad}}=\{u(\cdot)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(\alpha(\cdot),\eta(\cdot),\mu(\cdot),\nu(\cdot))\in BV([0,T];U_{\text{ad}})\,:\,TV(\alpha),TV(\eta),TV(\nu),TV(\mu)\leq M\}$
+for some universal constant $M$.
+That is, each individual control $\alpha,\eta,\mu,\nu$ is in $BV[0,T]$, the space of real-valued functions of bounded variation.
+To recall, this is the set of functions $z:[0,T]\to{\mathbb{R}}$ such that
+(3.4) $TV(z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\sup_{g\in C_{c}^{1}[0,T],\|g\|_{\infty}\leq 1}\int^{T}_{0}z(t)g^{\prime}(t)dt<\infty.$
+We remark on this choice of control functions after the proof of theorem 3.1. Our optimal control problem is then
 (3.5) $\min_{u\in\mathscr{U}_{\text{ad}}}J(x,u)\,\,\,\,\text{ subject to }\,\,\frac{dx}{dt}=f(x,u),\,\,\,\,x(0)=x_{0}$
 where $f(x,u)$ is defined by the right-hand side of (2.2) and $x_{0}\in{\mathbb{R}}^{6}$ has nonnegative components whose sum is bounded by $\tfrac{b}{\delta}$.
-Existence of an optimal control for this problem follows from classical optimal control theory.
+In this notation, we note that the unique a.e.
+solution $x$ of (2.2) corresponding to a control map $u\in\mathscr{U}_{ad}$ can be characterized as the unique function $x\in(L^{2}[0,T])^{6}$ satisfying the initial condition and
+(3.6) $\int^{T}_{0}\langle x(t),\phi^{\prime}(t)\rangle dt=\int^{T}_{0}\langle f(x(t),u(t)),\phi(t)\rangle dt,\,\,\,\,\,\forall\phi\in C^{1}_{c}([0,T];{\mathbb{R}}^{6}).$
+Because our admissible control set $U_{\text{ad}}$ is compact, $\mathscr{U}_{\text{ad}}\subset(L^{p}[0,T])^{4}$ for any $p\in[1,\infty)$ or $p=\infty.$ For simplicity, we prove the following existence theorem using the direct method considering convergence in the $L^{2}$ topology, though with small tweaks this argument would go through with other choices.
 Theorem 3.1.
-Assuming $c_{3},c_{4},c_{5},c_{6}$ are positive and $x_{0}$ has nonnegative components, there exists an optimal control $u\in\mathscr{U}_{\text{ad}}$ which solves the minimization problem (3.5).
-We omit the proof.
-Some key features of our system are that solutions of our system remain bounded and nonnegative as long as $x_{0}$ is componentwise nonnegative, $f(x,u)$ is smooth and affine in the control variables, and $r(x,u)$ is smooth and uniformly strongly convex in $u$ (when each of $c_{3},c_{4},c_{5},c_{6}$ are positive), and the control set $U_{\text{ad}}$ is convex and compact.
-These suffice to prove existence of an optimal control using [45, Theorem 2.1].
-In fact the particular dependence of running cost (3.1) on the state variables $x=(S,I,R,S^{*},I^{*},R^{*})$ could be significantly altered, as long as the smoothness in all variables and uniform strong convexity in control variables are maintained.
-We choose this particular running cost only to represent the concerns we imagine a policy-maker may have; other decisions are certainly valid.
-A necessary condition for optimality is provided by the Pontryagin minimum principle (PMP).
+Assuming $r$ and $\ell$ are proper, convex, and lower-semicontinuous and $x_{0}$ has nonnegative components, there exists an optimal control $u^{*}\in\mathscr{U}_{\text{ad}}$ and corresponding trajectory $x^{*}$ which solve the minimization problem (3.5).
+Proof.
+By proposition 2.1, there is a well-defined control-to-state operator $\mathcal{S}:\mathscr{U}_{\text{ad}}\to(H^{1}[0,T])^{6}$ so that for $u\in\mathscr{U}_{\text{ad}}$, $\mathcal{S}(u)=x\in(H^{1}[0,T])^{6}$ is the corresponding solution of (2.2).
+Note that the basic existence theory will only guarantee that $x\in(L^{2}[0,T])^{6}$, but then by the population bound (2.3), the right hand side of each equation in (2.2) defines a bounded, and hence $L^{2}[0,T]$, function so $x\in(H^{1}[0,T])^{6}$ by a bootstrapping argument.
+Furthermore $\mathcal{S}(u)=x$ remains uniformly bounded in $(H^{1}[0,T])^{6}$ among choices of $u\in\mathscr{U}_{\text{ad}}.$ We segment the proof into a few steps.
+In all following statements, we assume $\mathscr{U}_{\text{ad}}$ and $(H^{1}[0,T])^{6}$ are given the $L^{2}$ topology.
+Step 1: $\mathcal{S}$ is weakly closed.
+Suppose $u^{(k)}\in\mathscr{U}_{\text{ad}}$ is a sequence such that $u^{(k)}\rightharpoonup u$ in $(L^{2}[0,T])^{4}$ and $\mathcal{S}(u^{(k)})\rightharpoonup x$ in $(L^{2}[0,T])^{6}.$ Since $\mathscr{U}_{\text{ad}}$ is convex, it suffices to prove that it is strongly closed in $L^{2}$, and this will imply that it is weakly closed [15, Ch.
+5.1].
+That $\mathscr{U}_{\text{ad}}$ is closed under strong $L^{2}$ limits follows because if $z^{(k)}\to z$ strongly in $L^{2}$, then Hölder’s inequality gives
+$\left|\int^{T}_{0}z^{(k)}(t)g^{\prime}(t)dt-\int^{T}_{0}z(t)g^{\prime}(t)dt\right|\leq\|z^{(k)}-z\|_{L^{2}[0,T]}\|g^{\prime}\|_{L^{2}[0,T]}\to 0$
+for any $g\in C_{c}^{1}[0,T]$, and this proves that $TV(z^{(k)})\to TV(z)$ for total variation as defined in (3.4).
+In particular, strong $L^{2}$ limits respect bounds on total variation.
+Hence $\mathscr{U}_{\text{ad}}$ is weakly closed and so $u\in\mathscr{U}_{\text{ad}}$.
+Further, since $\{u^{(k)}\}$ is of uniformly bounded total variation, Helly’s selection theorem [25, Thm.
+5.5] allows us to pass to a subsequence along which $u^{(k)}\to u$ pointwise a.e.
+in $[0,T].$ (We do so without renaming the sequence, as we do whenever we mention a subsequence in the ensuing work.)
+Now $\mathcal{S}(u^{(k)})$ is a bounded sequence in the Hilbert space $(H^{1}[0,T])^{6}$, so it has subsequence converging weakly to some $\overline{x}\in(H^{1}[0,T])^{6}$.
+By the compact embedding $H^{1}[0,T]\hookrightarrow L^{2}[0,T]$ [24, Ch.
+5.7], passing to a further subsequence, we can ensure strong convergence to $\overline{x}$ in the $(L^{2}[0,T])^{6}$ topology.
+By uniqueness of limits, we must have $\overline{x}=x$, and thus passing to an even further subsequence, we have $\mathcal{S}(u^{(k)})\to x$ pointwise a.e.
+Finally, for each $k$, using the weak formulation (3.6), we have
+(3.7) $\int^{T}_{0}\langle[\mathcal{S}(u^{(k)})](t),\phi^{\prime}(t)\rangle dt=\int^{T}_{0}\langle f([\mathcal{S}(u^{(k)})](t),u^{(k)}(t)),\phi(t)\rangle dt,\,\,\,\,\,\forall\phi\in C^{1}_{c}([0,T];{\mathbb{R}}^{6}).$
+As $k\to\infty$, using strong $L^{2}$ convergence on the left hand side of (3.7), and pointwise a.e. convergence plus the dominated convergence theorem on the right hand side of (3.7), we see
+$\int^{T}_{0}\langle x(t),\phi^{\prime}(t)\rangle dt=\int^{T}_{0}\langle f(x(t),u(t)),\phi(t)\rangle dt,\,\,\,\,\,\forall\phi\in C^{1}_{c}([0,T];{\mathbb{R}}^{6}),$
+showing that $x$ is the solution of (2.2) corresponding to $u$. That is $\mathcal{S}(u)=x$, and $\mathcal{S}$ is weakly closed, which completes Step 1.
+Step 2: $\mathcal{S}$ is weak-strong continuous.
+Suppose $u^{(k)}\rightharpoonup u$ weakly in $L^{2}$.
+We need to prove that $\mathcal{S}(u^{(k)})\to\mathcal{S}(u)$ in $L^{2}.$ Once again, $\{\mathcal{S}(u^{(k)}\}$ is a uniformly bounded sequence in $(H^{1}[0,T])^{6}$, and so any subsequence $\{\mathcal{S}(u^{(k_{\ell})}\}$ is still uniformly bounded $(H^{1}[0,T])^{6}$.
+By the compact embedding $H^{1}[0,T]\hookrightarrow L^{2}[0,T]$, $\{\mathcal{S}(u^{(k_{\ell})}\}$ has a further subsequence $\{S(u^{(k_{\ell_{m}})}\}$ converging strongly (hence also weakly) to some function $x\in(L^{2}[0,T])^{6}$.
+By weak closure, we must have $\mathcal{S}(u)=x$, so we have proven that every subsequence of $\{\mathcal{S}(u^{(k)}\}$ has a further subsequence converging strongly to $\mathcal{S}(u)$ in $(L^{2}[0,T])^{6}$.
+This implies that $\mathcal{S}(u^{(k)})\to\mathcal{S}(u)$ strongly in $(L^{2}[0,T])^{6}$ by Urysohn’s subsequence principle [53, Ch.
+2.1.17].
+Step 3: The cost functional is lower-semicontinuous so (3.5) has a solution.
+Because the control-to-state operator is well-defined, (3.5) is equivalent to the reduced problem
+(3.8) $\min_{u\in\mathscr{U}_{\text{ad}}}\mathcal{J}(u)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=J(\mathcal{S}(u),u).$
+Since $r,\ell$ are proper, convex, and lower-semicontinuous, they are bounded below, and since $u,\mathcal{S}(u)$ are bound bounded functions, this implies $\mathcal{J}(u)$ is bounded from below and thus has a finite infimum $\mathcal{J}_{0}$.
+Let $\{u^{(k)}\}$ be a minimizing sequence of controls, so that $\mathcal{J}(u^{(k)})\to\mathcal{J}_{0}.$ Since $\{u^{(k)}\}$ is a bounded sequence in the Hilbert space $(L^{2}[0,T])^{4}$, we can pass to a subsequence that converges weakly to some $u\in\mathscr{U}_{\text{ad}}$.
+By weak closure and weak-strong continuity, $\mathcal{S}(u^{(k)})\to x\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathcal{S}(u)$ strongly in $(L^{2}[0,T])^{6}.$ Since $r,\ell$ are proper, convex, and lower-semicontinuous, the maps
+$x\mapsto\int^{T}_{0}r(x(t))dt,\,\,\,\,\,\,\,\,u\mapsto\int^{T}_{0}\ell(u(t))dt$
+are weakly lower-semicontinuous, and thus
+$\mathcal{J}_{0}\leq\mathcal{J}(u)=J(x,u)$ $=\int^{T}_{0}(r(x(t))+\ell(u(t)))dt$
+$\leq\liminf_{k\to\infty}\int^{T}_{0}\Big((r([\mathcal{S}(u^{(k)})](t))+\ell(u^{(k)}(t))\Big)dt=\liminf_{k\to\infty}\mathcal{J}(u^{(k)})=J_{0}.$
+Thus $u\in U_{\text{ad}}$ and the associated solution $x$ of (2.2) are a solution of (3.5). ∎
+We prove one further regularity result for the control-to-state map. While this results was not necessary in the proof above, it will have bearing in section 4.
+Lemma 3.2.
+The control-to-state operator $\mathcal{S}:\mathscr{U}_{\text{ad}}\to(H^{1}[0,T])^{6}$ is Lipschitz continuous when both the source and target spaces are given the $L^{2}$ topology.
+Proof.
+Let $u,v\in\mathscr{U}_{\text{ad}}$ be two control maps, and let $x=\mathcal{S}(u)$ and $y=\mathcal{S}(v)$ be the associated solutions of (2.2). Then for any $t\in[0,T],$
+$\left|x(t)-y(t)\right|_{2}\leq\int^{t}_{0}\left|f(x(s),u(s))-f(y(s),v(s))\right|_{2}ds.$
+By proposition 2.1, solutions of (2.2) are uniformly bounded over choices of controls, and the controls themselves are uniformly bounded.
+Thus since $f(x,u)$ is smooth in both $x$ and $u$, it is, in particular, Lipschitz continuous, so
+$\left|x(t)-y(t)\right|_{2}\leq L\int^{t}_{0}\Big(\left|x(s)-y(s)\right|_{2}+\left|u(s)-v(s)\right|_{2}\Big)ds.$
+Thus Grönwall’s inequality gives
+$\left|x(t)-y(t)\right|_{2}\leq Le^{T}\int^{t}_{0}\left|u(s)-v(s)\right|_{2}ds\leq Le^{T}\int^{T}_{0}\left|u(s)-v(s)\right|_{2}ds.$
+Squaring this and invoking Jensen’s inequality yields
+$\left|x(t)-y(t)\right|^{2}_{2}\leq L^{2}Te^{2T}\|u-v\|^{2}_{(L^{2}[0,T])^{4}}$
+and integrating on $[0,T]$ shows that
+$\|x-y\|^{2}_{(L^{2}[0,T])^{6}}\leq(LTe^{T})^{2}\|u-v\|^{2}_{(L^{2}[0,T])^{4}}$
+proving the claim. ∎
+Remark.
+A remark is in order regarding the specific choice of our control set $\mathscr{U}_{\text{ad}}$.
+If the dynamics are control affine, it is enough to assume $u\in(L^{2}[0,T])^{4}$.
+However, our dynamics are not control affine.
+Thus in Step 1 of the proof, some higher regularity is required to ensure a stronger notion of convergence of $u^{(k)}\to u$, and allow us to pass to limit in (3.7).
+The proof above would go through a bit more easily if we assume even more regularity–for example $u\in(H^{1}[0,T])^{4}$–but this is highly undesirable, because it would preclude the possibility of discontinuous control maps.
+Discontinuous controls arise naturally when enforcing $L^{1}$ control costs since these commonly lead to bang-bang control strategies (see [33, Ch.
+17] or [52, 57]).
+The assumption that the controls have uniformly bounded total variation allows for the use of Helly’s selection theorem [25, Thm.
+5.5] to pass to subsequence which converges pointwise a.e., while still maintaining strong closure of $\mathscr{U}_{\text{ad}}$ in the $L^{2}$ topology, and also allowing for discontinuous control maps.
+Given existence of optimal controls, a necessary condition for optimality is provided by the Pontryagin minimum principle (PMP).
 Using state variables $x=(S,I,R,S^{*},I^{*},R^{*})$, costate variables $p=(p_{S},p_{I},p_{R},p_{S^{*}},p_{I^{*}},p_{R^{*}})$, and control variables $u=(\alpha,\eta,\mu,\nu),$ we define the Hamiltonian
-(3.6) $\begin{split}H(x,p,u)&=\langle p,f(x,u)\rangle+r(x,u),\end{split}$
-where $r(x,u)$ is the running cost (3.1) and $f(x,u)$ is defined by the right-hand side of (2.2).
+(3.9) $\begin{split}\mathcal{H}(x,p,u)&=\langle p,f(x,u)\rangle+r(x)+\ell(u),\end{split}$
+where $r,\ell$ are as in (3.1) and $f(x,u)$ is defined by the right-hand side of (2.2).
 Then the PMP states that if $x(\cdot),u(\cdot)$ are an optimal state and control pair for (3.5), then there is a corresponding costate trajectory $p(\cdot)$ such that
@@
 (3.12) $u(t)$ $=\operatorname*{argmin}_{\hat{u}\in U_{\text{ad}}}\mathcal{H}(x(t),p(t),\hat{u}),\,\,\,\,\,\text{ for all }\,\,\,t\in[0,T].$
-We direct the reader to several classical sources for proofs and discussions regarding the PMP; for example, [2, §5], [29, §4], [42, §1,2], [49, §12].
+We direct the reader to several classical sources for proofs and discussions regarding the PMP; for example, [3, §5], [34, §4], [48, §1,2], [58, §12].
 We note that (3.10) is exactly the original dynamics (2.2).
@@
 The condition $p(T)=0$ in (3.11) is called the transversality condition (discussed in the aforementioned sources), and takes this simplified form because we consider a fixed time, free end-point problem with no terminal cost.
-We will use the forward-backward sweep method described in [28, §4] to numerically solve (3.7)-(3.9).
-We fully detail the algorithm shortly, but first, to make all computations entirely explicit, we write out (3.8) and (3.9) for our specific choice of cost functional (3.2).
-In our case, the Hamiltonian can be written
-(3.10) $\begin{split}\mathcal{H}(x,p,u)&=p_{S}\Big{(}(1-\xi)b-\beta(1-\alpha)S(I+I^{*})-(\overline{\mu}-\mu)SN^{*}+\nu S^{*}-\delta S\Big{)}\\ &\hskip 28.45274pt+p_{I}\Big{(}\beta(1-\alpha)S(I+I^{*})-(\gamma+\eta)I-(\overline{\mu}-\mu)IN^{*}+\nu I^{*}-\delta I\Big{)}\\ &\hskip 28.45274pt+p_{R}\Big{(}(\eta+\gamma)I-(\overline{\mu}-\mu)RN^{*}+\nu R^{*}-\delta R\Big{)}\\ &\hskip 28.45274pt+p_{S^{*}}\Big{(}\xi b-\beta S^{*}(I+I^{*})+(\overline{\mu}-\mu)SN^{*}-\nu S^{*}-\delta S^{*}\Big{)}\\ &\hskip 28.45274pt+p_{I^{*}}\Big{(}\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu)IN^{*}-\nu I^{*}-\delta I^{*}\Big{)}\\ &\hskip 28.45274pt+p_{R^{*}}\Big{(}\gamma I^{*}+(\overline{\mu}-\mu)RN^{*}-\nu R^{*}-\delta R^{*}\Big{)}\\ &\hskip 28.45274pt+c_{1}(I+I^{*})+c_{2}N^{*}+c_{3}\frac{\alpha^{2}}{2}+c_{4}\frac{\eta^{2}}{2}+c_{5}\frac{\mu^{2}}{2}+c_{6}\frac{\nu^{2}}{2}.\end{split}$
-Thus, (3.8) reads
-(3.11) $\begin{split}\frac{dp_{S}}{dt}&=\beta(1-\alpha)(I+I^{*})(p_{S}-p_{I})+(\overline{\mu}-\mu)N^{*}(p_{S}-p_{S^{*}})+\delta p_{S},\\ \frac{dp_{I}}{dt}&=\beta(1-\alpha)S(p_{S}-p_{I})+\beta S^{*}(p_{S^{*}}-p_{I^{*}})+(\gamma+\eta)(p_{I}-p_{R})+\delta p_{I}-c_{1}\\ &\hskip 28.45274pt+(\overline{\mu}-\mu)N^{*}(p_{I}-p_{I^{*}}),\\ \frac{dp_{R}}{dt}&=(\overline{\mu}-\mu)N^{*}(p_{R}-p_{R^{*}})+\delta p_{R},\\ \frac{dp_{S^{*}}}{dt}&=\beta(I+I^{*})(p_{S^{*}}-p_{I^{*}})+\nu(p_{S^{*}}-p_{S})+\delta p_{S^{*}}-c_{2}\\ &\hskip 28.45274pt+(\overline{\mu}-\mu)\big{(}S(p_{S}-p_{S^{*}})+I(p_{S}-p_{S^{*}})+R(p_{R}-p_{R^{*}})\big{)},\\ \frac{dp_{I^{*}}}{dt}&=\beta(1-\alpha)S(p_{S}-p_{I})+\beta S^{*}(p_{S^{*}}-p_{I^{*}})+\gamma(p_{I^{*}}-p_{R^{*}})+\nu(p_{I^{*}}-p_{I})-c_{1}-c_{2}\\ &\hskip 28.45274pt+(\overline{\mu}-\mu)\big{(}S(p_{S}-p_{S^{*}})+I(p_{S}-p_{S^{*}})+R(p_{R}-p_{R^{*}})\big{)}+\delta p_{I^{*}},\\ \frac{dp_{R^{*}}}{dt}&=\nu(p_{R^{*}}-p_{R})+(\overline{\mu}-\mu)\big{(}S(p_{S}-p_{S^{*}})+I(p_{S}-p_{S^{*}})+R(p_{R}-p_{R^{*}})\big{)}+\delta p_{R^{*}}-c_{2},\end{split}$
-along with zero terminal condition $p_{X}(T)=0$ for each $X\in\{S,I,R,S^{*},I^{*},R^{*}\}.$
-Because the Hamiltonian is quadratic in the control variables $u=(\alpha,\eta,\mu,\nu)$, one can easily resolve the optimality condition (3.9) by simply minimizing a quadratic function and projecting back onto the admissible set of control values.
-Suppressing dependence on $t$, once the state and costate variables are known, optimal controls are given by
-(3.12) $\begin{split}\alpha&=\min\left\{\max\left\{\frac{\beta S(I+I^{*})(p_{I}-p_{S})}{c_{3}},\,0\right\},\,1\right\},\\ \eta&=\min\left\{\max\left\{\frac{I(p_{I}-p_{R})}{c_{4}},\,0\right\},\,\overline{\eta}\right\},\\ \mu&=\min\left\{\max\left\{\frac{[S(p_{S^{*}}-p_{S})+I(p_{I^{*}}-p_{I})+R(p_{R^{*}}-p_{R})]N^{*}}{c_{5}},\,0\right\},\,\overline{\mu}\right\},\\ \nu&=\min\left\{\max\left\{\frac{S^{*}(p_{S^{*}}-p_{S})+I^{*}(p_{I^{*}}-p_{I})+R^{*}(p_{R^{*}}-p_{R})}{c_{6}},\,0\right\},\,\overline{\nu}\right\}.\end{split}$
-Given this, the forward-backward sweep method is detailed in algorithm 1.
-We include some basic implementation notes.
-For the initial guess, we always use $u^{(0)}\equiv 0$.
-Step (iv) in the iteration is a relaxation step which may not be necessary but, empirically, accelerates convergence.
-For convergence criterion, we halt the iteration when the relative change in control variables (measured in supremum norm) is below some specified tolerance.
-Using the suggestion of [28, §4], we rearrange the condition slightly so as to avoid division by zero in the case that zero control is optimal.
-Thus our convergence criterion is
-(3.13) $(\text{tol})\cdot\|\alpha^{(k+1)}\|_{\infty}-\|\alpha^{(k+1)}-\alpha^{(k)}\|_{\infty}\geq 0,$
-where “tol” is the convergence tolerance.
-While (3.13) is written for $\alpha$, we enforce this criterion for each of $\alpha,\eta,\mu,\nu$ and only halt the iteration once all of them satisfy the condition.
-Algorithm 1 Forward-Backward Sweep Method for Resolving Optimal Controls
-Input: initial condition $x_{0}=(S_{0},I_{0},R_{0},S^{*}_{0},I^{*}_{0},R^{*}_{0})\in{\mathbb{R}}^{6}_{+}$, parameter values $b,\delta,\beta,\gamma,\overline{\eta},\xi,\overline{\mu},\overline{\nu}$, cost weights $c_{1},\ldots,c_{6}$, a relaxation parameter $\kappa\in[0,1]$ and an initial guess for the control map $u^{(0)}=(\alpha^{(0)},\eta^{(0)},\mu^{(0)},\nu^{(0)})$.
-Set $k=0$ and do the following:
+In the ensuing section, we describe numerically resolving optimal controls via the sequential quadratic Hamiltonian method [10].
+4. Numerical Methods
+In this section, we consider the numerical resolution of optimal control maps via the sequential quadratic Hamiltonian (SQH) method [10].
+Because this is a fairly new numerical method, and to make this manuscript as self-contained as possible, we discuss some basic facets of the method and prove convergence of the iteratively defined control maps.
+A full discussion of the method with worked examples can be found in [10].
+The SQH method is an iterative method for resolving control maps in a variety of scenarios involving optimal control of ordinary or partial differential equations.
+The basic strategy is to successively update the approximation of the control maps using quadratic perturbations to the true Hamiltonian (3.9).
+Specifically, for $\varepsilon>0$, defined the auxiliary Hamiltonian
+(4.1) $\mathcal{H}_{\varepsilon}(x,p,u,\overline{u})\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\mathcal{H}(x,p,\overline{u})+\varepsilon\left|u-\overline{u}\right|^{2}_{2}.$
+The quadratic term above introduces relaxation, so that solving
+$u^{\text{new}}(t)=\operatorname*{argmin}_{w\in U_{\text{ad}}}\mathcal{H}_{\varepsilon}(x(t),p(t),u(t),w)$
+results in $u^{\text{new}}$ which balances minimizing the Hamiltonian with a penalty for straying too far from the currently employed control strategy $u$.
+Given this definition, the full SQH algorithm is detailed in algorithm 1.
+The basic strategy of the algorithm is to successively update $x,p$ and $u$ by solving the state and co-state equations and resolving controls via (4.1).
+One significant novelty is the inclusion of step (v) in the iteration, whereupon one checks the condition
+(4.2) $J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)\leq-\rho\,\tau.$
+If this condition is met, it is determined that the new controls $u^{(k+1)}$ have achieved a sufficient decrease in the cost functional.
+In this case, the new controls $u^{(k+1)}$ are accepted, $\varepsilon$ is decreased by setting $\varepsilon\leftarrow\zeta\varepsilon$ for some $\zeta\in(0,1)$, and the iteration continues with the updated auxiliary Hamiltonian.
+If the condition is not met, there was not sufficient decrease in the cost functional, so $\varepsilon$ is increased by setting $\varepsilon\leftarrow\lambda\varepsilon$ for some $\lambda>1$ and $u^{(k+1)}$ is recomputed before continuing.
+Here $\lambda,\zeta$ are tunable parameters, as is $\rho$ in (4.2), which determines the level to which $u^{(k+1)}$ must decrease the cost functional to be considered acceptable.
+Algorithm 1 Sequential Quadratic Hamiltonian Method for Resolving Optimal Control Maps
+(1) Input: initial approximation $u^{0}$, maximum number of iterations $k_{max}$, tolerance $\kappa>0$, and hyperparamaters $\varepsilon>0$, $\lambda>1$, $\rho>0$, and $\zeta\in\left(0,1\right)$.
+(2) Set $k=0$ and compute the solution $x^{(0)}=(S^{(0)},I^{(0)},R^{(0)},S^{*,(0)},I^{*,(0)},R^{*,(0)})$ to (2.2) with control $u=u^{(0)}$.
+(3) Perform the following iteration.
 repeat
-  (i) Resolve $x^{(k+1)}=(S^{(k+1)},I^{(k+1)},R^{(k+1)},(S^{*})^{(k+1)},(I^{*})^{(k+1)},(R^{*})^{(k+1)})$ as the solution of (2.2) with initial condition $x_{0}$, using control variables $u^{(k)}$
-  (ii) Resolve $p^{(k+1)}=(p_{S}^{(k+1)},p_{I}^{(k+1)},p_{R}^{(k+1)},p_{S^{*}}^{(k+1)},p_{I^{*}}^{(k+1)},p_{R^{*}}^{(k+1)})$ as the solution of (3.11) with zero terminal condition, using control variables $u^{(k)}$ and state variables $x^{(k+1)}$
-  (iii) Define $u^{*}=(\alpha^{(k+1)},\eta^{(k+1)},\mu^{(k+1)},\nu^{(k+1)})$ from (3.12) using state variables $x^{(k+1)}$ and costate variables $p^{(k+1)}$
-  (iv) Define $u^{(k+1)}=(1-\kappa)u^{(k)}+\kappa u^{*}$ and increment $k$
-until convergence criteria are met
-return $u=u^{(k)}$, $x=x^{(k)}$, $p=p^{(k)}$: respectively the optimal control, state, and costate trajectories
+  (i) Compute the costate dynamics $p^{(k)}$ obeying (3.11) with $x=x^{(k)}$ and $u=u^{(k)}$.
+  (ii) Compute $u^{(k+1)}$ via:
+$u^{(k+1)}(t)=\operatorname*{argmin}_{w\in U_{\text{ad}}}\,\mathcal{H}_{\varepsilon}\left(x^{(k)}(t),p^{(k)}(t),u^{(k)}(t),w\right),\,\,\,\,\,\,\,t\in[0,T].$
+  (iii) Compute the solution $x^{(k+1)}$ to the (2.2) with control $u=u^{(k+1)}.$
+  (iv) Compute $\tau=\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}$.
+  (v) if: $\left\{J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)\leq-\rho\,\tau\right\}$ then set $\varepsilon\leftarrow\zeta\varepsilon$ and continue to step (vi),
+     else if: $\left\{J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)>-\rho\,\tau\right\}$ then set $\varepsilon\leftarrow\lambda\varepsilon$ and go to step (ii).
+  (vi) Set $k\leftarrow k+1.$
+until $k=k_{\text{max}}$ or $\tau<\kappa$
+return $(x^{*},u^{*})$ - the most recently computed state and control trajectories
+We prove a basic convergence result for algorithm 1, after establishing a lemma.
+Lemma 4.1.
+Assume that the state cost function $r$ from (3.1) has a Lipschitz continuous gradient.
+Let $u,v\in\mathscr{U}_{\text{ad}}$ be two control maps with associated solutions $x=\mathcal{S}(u)$ and $y=\mathcal{S}(v)$ of (2.2).
+Then there is a constant $c>0$ such that
+(4.3) $J(x,u)-J(y,v)\leq\int^{T}_{0}\Big(\mathcal{H}(x(t),p(t),u(t))-\mathcal{H}(x(t),p(t),v(t))\Big)dt+c\|u-v\|^{2}_{(L^{2}[0,T])^{4}}$
+where $p$ is the co-state trajectory corresponding to $(x,u)$.
+Proof.
+We compute, dropping depending on $t$ for brevity,
+(4.4) $J(x,u)-J(y,v)$ $=\int^{T}_{0}\Big(r(x)+\ell(u)-r(y)-\ell(v)\Big)dt$
+$=\int^{T}_{0}\Big([\mathcal{H}(x,p,u)-\mathcal{H}(y,p,v)]+\left\langle p,f(y,v)-f(x,u)\right\rangle\Big)dt$
+$=\int^{T}_{0}[\mathcal{H}(x,p,u)-\mathcal{H}(x,p,v)]dt$
+$\hskip 56.9055pt+\int^{T}_{0}\Big([\mathcal{H}(x,p,v)-\mathcal{H}(y,p,v)]+\left\langle p,\tfrac{d}{dt}(y-x)\right\rangle\Big)dt.$
+In the last equality of (4.4), the first term on the right hand side is as it appears in (4.3), so we work on the last line, which we call $I$. We see
+(4.5) $I$ $=\int^{T}_{0}\Big([\mathcal{H}(x,p,v)-\mathcal{H}(y,p,v)]+\left\langle p,\tfrac{d}{dt}(y-x)\right\rangle\Big)dt$
+$=\int^{T}_{0}\left(\int^{1}_{0}-\left\langle\nabla_{x}\mathcal{H}(x+\lambda(y-x),p,v),y-x\right\rangle d\lambda+\left\langle p,\tfrac{d}{dt}(y-x)\right\rangle\right)dt.$
+Now integrating by parts to pass the derivative onto $p$, the boundary terms vanish since $x(0)=y(0)$ and $p(T)=0$. So using (3.11), (4.5) becomes
+(4.6) $I=\int^{T}_{0}\int^{1}_{0}\left\langle\nabla_{x}\mathcal{H}(x,p,u)-\nabla_{x}H(x+\lambda(y-x),p,v),y-x\right\rangle d\lambda dt.$
+Now $\nabla_{x}\mathcal{H}=(J_{f}(x,u))^{T}p+\nabla r$ where $J_{f}$ is the Jacobian matrix of $f$.
+Since $f$ is smooth (and its arguments remain bounded) and since $\nabla r$ is Lipschitz by assumption, $\nabla_{x}\mathcal{H}$ is Lipschitz.
+Thus (4.6) implies
+(4.7) $I$ $\leq L\int^{T}_{0}\int^{1}_{0}(\lambda\left|x-y\right|^{2}_{2}+\left|u-v\right|_{2}\left|x-y\right|_{2})d\lambda dt$
+$\leq\frac{L}{2}\|x-y\|^{2}_{(L^{2}[0,T])^{6}}+L\|x-y\|_{(L^{2}[0,T])^{6}}\|u-v\|_{(L^{2}[0,T])^{4}}.$
+Finally, lemma 3.2 tells us that the control-to-state map is Lipschitz, whereupon (4.7) yields
+$I\leq c\|u-v\|^{2}_{(L^{2}[0,T])^{2}}$
+and inserting this last bound into (4.4) concludes the proof. ∎
+Theorem 4.2.
+Under the assumptions of lemma 4.1, let $\left(x^{(k)},u^{(k)}\right)$ and $\left(x^{(k+1)},u^{(k+1)}\right)$ be successive iterates generated by the SQH method (algorithm 1).
+Then there is $c>0$ such that
+(4.8) $J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)\leq-(\varepsilon-c)\,\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}.$
+for the current value of $\varepsilon>0$ in algorithm 1. In particular, this implies
+(4.9) $J\left(x^{(k+1)},u^{(k+1)}\right)-J\left(x^{(k)},u^{(k)}\right)\leq-\rho\,\tau$
+for $\varepsilon\geq\rho+c$ and $\tau=\|u^{k+1}-u^{k}\|^{2}_{(L^{2}[0,T])^{3}}$.
+Proof.
+In step (ii), we resolve $u^{(k+1)}$ such that
+$\mathcal{H}_{\varepsilon}\left(x^{(k)}(t),p^{(k)}(t),u^{(k)}(t),u^{(k+1)}(t)\right)\leq\mathcal{H}_{\varepsilon}\left(x^{(k)}(t),p^{(k)}(t),u^{(k)}(t),w\right),$
+for all $w\in U_{\text{ad}}\text{ and }t\in[0,T].$ In particular, choosing $w=u^{(k)}(t)$ at each $t\in[0,T]$ (and then dropping the argument $t$ for brevity), we see
+$\mathcal{H}_{\varepsilon}\left(x^{(k)},p^{(k)},u^{(k)},u^{(k+1)}\right)\leq\mathcal{H}_{\varepsilon}\left(x^{(k)},p^{(k)},u^{(k)},u^{(k)}\right)=\mathcal{H}\Big(x^{(k)},p^{(k)},u^{(k)}\Big).$
+Inserting the definition of $\mathcal{H}_{\varepsilon}$, this yields
+$\mathcal{H}\Big(x^{(k)},p^{(k)},u^{(k+1)}\Big)-\mathcal{H}\Big(x^{(k)},p^{(k)},u^{(k)}\Big)\leq-\varepsilon\big|u^{(k+1)}-u^{(k)}\big|^{2}_{2}.$
+Then from lemma 4.1, we have
+$J\Big(x^{(k+1)},u^{(k+1)}\Big)-J\Big(x^{(k)},u^{(k)}\Big)$ $\leq\int^{T}_{0}\Big(H\Big(x^{(k)},p^{(k)},u^{(k+1)}\Big)-\mathcal{H}\Big(x^{(k)},p^{(k)},u^{(k)}\Big)\Big)dt$
+$\hskip 85.35826pt+c\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}$
+$\leq-(\varepsilon-c)\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}$
+as desired. ∎
+Recall, if sufficient decrease of the cost functional is not achieved, $\varepsilon$ is successively increased.
+Theorem 4.2 tells us that once $\varepsilon\geq\rho+c$, we are guaranteed to achieve sufficient decrease of the cost functional.
+Given this, we prove that the numerical iterates do indeed converge.
+Theorem 4.3.
+The sequence of iterates $\{(x^{(k)},u^{(k)})\}$ created by the SQH algorithm satisfy the following.
+  • (a)
+    $\{J\Big(x^{(k)},u^{(k)}\Big)\}$ decreases in $k$, and thus converges to some $J^{*}\geq\min_{u\in\mathscr{U}_{\text{ad}}}J(x,u)$
+  • (b)
+    $\|u^{(k+1)}-u^{(k)}\|_{(L^{2}[0,T])^{4}},\,\|x^{(k+1)}-x^{(k)}\|_{(L^{2}[0,T])^{6}}\to 0$ as $k\to\infty$
+Proof.
+Monotonicity of $J(x^{(k)},u^{(k)})$ follows directly from (4.9) following the observation above that eventually $\varepsilon$ will be increased enough that $\varepsilon\geq\rho+c$.
+Convergence then follows by the monotone convergence theorem, so (a) is proven.
+For (b), we rearrange (4.9) to see that when $u^{(k+1)}$ is resolved,
+$\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}\leq\frac{1}{\rho}\Big(J\Big(x^{(k)},u^{(k)}\Big)-J\Big(x^{(k+1)},u^{(k+1)}\Big)\Big).$
+Thus we have the telescoping sum
+$\sum^{K}_{k=0}\|u^{(k+1)}-u^{(k)}\|^{2}_{(L^{2}[0,T])^{4}}\leq\frac{1}{\rho}\Big(J\Big(x^{(0)},u^{(0)}\Big)-J\Big(x^{(K+1)},u^{(K+1)}\Big)\Big).$
+As $K\to\infty$ the right hand side above converges, so this shows that the left hand side is a convergent series of positive terms, implying that the summands go to zero.
+That $\|x^{(k+1)}-x^{(k)}\|_{(L^{2}[0,T])^{6}}\to 0$ as well follows by the Lipschitz continuity of the control-to-state map proven in lemma 3.2.
+∎
+With this, we move on to simulations and discussion wherein we define a cost functional discuss the optimal control strategies that result in the case of different policy-maker preferences.
 5. Results & Discussion
-We implement algorithm 1 in MATLAB and demonstrate the results in several different cases.
-Specifically, we are interested in observing the differing behavior of optimal control maps in response to changes in the policy-maker’s preferences, expressed via the weights $c_{1},\ldots,c_{6}$ from (3.1).
-For all simulations, we use relaxation parameter $\kappa=0.8$ and a time horizon of $T=100$, and for the sake of numerically resolving the solutions of (2.2) and (3.11), we use an explicit Euler scheme with uniform discretization of $[0,T]$ with step size $\Delta t=0.1$, and a convergence tolerance of $\text{tol}=10^{-3}.$ The code which generated all of these results is available online.
-[^1] https://github.com/chparkinson/Opt_Control_SIR_With_Noncompliance_As_Social_Contagion
-What remains is to specify initial conditions and the parameters $b,\delta,\beta,\gamma,\overline{\eta},\xi,\overline{\mu},\overline{\nu}$.
+We implement algorithm 1 in MATLAB and demonstrate the results in several different cases. Specifically, we consider the cost functional (3.1) with
+(5.1) $r(x)$ $=c_{1}(I+I^{*})+c_{2}(S^{*}+I^{*}+R^{*}),$
+$\ell(u)$ $=d_{1}(\left|\alpha\right|+\left|\eta\right|+\left|\mu\right|+\left|\nu\right|)+\frac{d_{2}}{2}(\alpha^{2}+\eta^{2}+\mu^{2}+\nu^{2}).$
+where $c_{1},c_{2},d_{1},d_{2}$ are nonnegative weights that express the policy-maker’s preferences.
+We are interested in observing the differing behavior of optimal control maps in response to changes in the policy-maker’s preferences.
+That is, in each scenario below, we fix these weights and use the SQH method to resolve approximate optimal control maps.
+For all simulations we set $T=100$, and for the sake of numerically resolving the solutions of (2.2) and (3.11), we use an explicit Euler scheme with uniform discretization of $[0,T]$ with step size $\Delta t=0.1$.
+What remains is to specify initial conditions and the parameters $b,\delta,\beta,\gamma,\overline{\eta},\xi,\mu^{*},\overline{\mu},\overline{\nu}$.
 To enhance readability, we include a description of each of these parameters and the values they take in each scenario.
@@
 Before displaying results, we briefly discuss the bearing of the controls.
-Recall that the cost (3.2) for the policy-maker is split into three parts: (C1) cost associated with infections (weighted by the constant $c_{1}$), (C2) costs associated with noncompliance (weighted by the constant $c_{2}$) and (C3) cost associated with controls (weighted by constants $c_{3},\ldots,c_{6}$).
+Recall that the cost (3.1) for the policy-maker is split into three parts: (C1) cost associated with infections (weighted by the constant $c_{1}$), (C2) costs associated with noncompliance (weighted by the constant $c_{2}$) and (C3) cost associated with controls (weighted by constants $d_{1},d_{2}$).
 By proposition 2.2, any use of controls will help decrease cost (C1) by slowing the spread of infections.
@@
 We mention specifically that the absolute size of the weights is not important, but rather their size relative to each other.
-Accordingly, in much of the below we fix the control cost weights $c_{3},\ldots,c_{6}=1$ and vary the cost weights $c_{1},c_{2}$ associated with total infections and total amount of noncompliance, respectively.
-In In certain cases, we would like to “turn off” individual controls; to do so, one can simply set that control to zero (this could equivalently be accomplished by sending the control’s cost weight to infinity).
+Accordingly, in much of the below we fix the control cost weights $d_{1}=d_{2}=0.1$ and vary the cost weights $c_{1},c_{2}$ associated with total infections and total amount of noncompliance, respectively.
+In certain cases, we would like to “turn off” individual controls; to do so, one can simply set the upper bound for that control to zero (this could equivalently be accomplished by sending the control’s cost weight to infinity).
 Lastly, as baseline values for infection and noncompliance cost weights, we choose $c_{1}=1$ and $c_{2}=0.1$.
@@
 The reason that $c_{2}$ must be significantly smaller is that with no controls, noncompliance is a permanent state, whereas infection is not, which means that noncompliance has the potential to contribute much more to the total cost.
+Finally, we emphasize that for each separate plot of controls below, we are changing certain and resolving optimal controls again using the SQH method (i.e., none of these plots contain synthetically designed controls; all of the controls presents are solutions of an optimal control problem with different parameters and/or cost weights).
 Parameter Description Scen. 1 Scen. 2 Scen. 3
@@
 $\gamma$ disease recovery rate 0.2 0.2 0.2
+$\overline{\alpha}$ max. achievable decrease in mixing due to compliance 0.5 0.5 0.5
 $\overline{\eta}$ maximum achievable treatment efficacy 0.1 0.1 0.1
 $\xi$ noncompliant portion of new population 0 0.3 var.
-$\overline{\mu}$ baseline infection rate for noncompliance 0.1 0.2 0.5
+$\mu^{*}$ baseline infection rate for noncompliance 0.1 0.2 0.3
+$\overline{\mu}$ max. achievable decrease in inf. rate of noncomp. 0.05 0.1 0.15
 $\overline{\nu}$ maximum achievable rehab. rate for noncompliance 0.1 0.1 0.1
 $c_{1}$ cost weight on infections var. var. 1
-$c_{2}$ cost weight on infections var. var. 0.1
-$c_{3},\ldots,c_{6}$ cost weights on controls 1 1 1
+$c_{2}$ cost weight on noncompliance var. 0.01 0.05
+$d_{1}$ $L^{1}$ cost weights on controls 0.1 0.1 0.1
+$d_{2}$ $L^{2}$ cost weight on controls var. 0.1 0.1
 Figure 1.
 Table of parameters for each scenario.
-Recall in particular that $\overline{\eta},\overline{\mu},\overline{\nu}$ are upper bounds for the control variables $\eta,\mu,\nu$, which represent increase in recovery rate due to seeking treatment, decrease in spread rate of noncompliance due to public health initiatives, and recovery rate for noncompliance due to educational campaigns, respectively.
+Recall in particular that $\overline{\eta},\mu^{*},\overline{\nu}$ are upper bounds for the control variables $\eta,\mu,\nu$, which represent increase in recovery rate due to seeking treatment, decrease in spread rate of noncompliance due to public health initiatives, and recovery rate for noncompliance due to educational campaigns, respectively.
 Not mentioned is the control variable $\alpha$, representing decrease in infection rate among the compliant susceptible population due to nonpharmaceutical intervention measures, since $\alpha$ will take values in $[0,1].$
@@
 With no controls, the reproductive ratio for this scenario is $\mathscr{R}_{0}=\tfrac{b}{\delta}\cdot\frac{\beta}{\gamma+\delta}=\frac{0.4}{0.21}\approx 1.9.$ We note in particular that $\xi=0$, meaning that all newly introduced members of the population are compliant, and we are in the regime of theorem 2.3(i),(ii).
-In this case, there are two DFEs, that, when controls are constant, take the forms $(s_{1},s^{*}_{1})=(\tfrac{b}{\delta},0)$ and $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$.
+In this case, there are two DFEs, that, when controls are constant, take the forms $(s_{1},s^{*}_{1})=(\tfrac{b}{\delta},0)$ and $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})$.
 If the controls are constant, the second is stable in the case that it is physically meaningful (i.e.
-$\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$); otherwise, the first is stable.
+$\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\mu^{*}-\mu}$); otherwise, the first is stable.
 Clearly to minimize (C1) and (C2), it behooves the policy-maker to drive the system toward $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$.
-If control cost is no issue, the policy-maker can increase $\mu,\nu$ to ensure that $(s_{1},s_{1}^{*})$ is stable, and then decrease $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\tfrac{\beta(1-\alpha)}{\gamma+\eta+\delta}$ using $\alpha,\mu$ so as to fall into the stability regime of theorem 2.3(i).
+If control cost is no issue, the policy-maker can increase $\mu,\nu$ to ensure that $(s_{1},s_{1}^{*})$ is stable, and then decrease $\mathscr{R}_{0}(\tfrac{b}{\delta},0)=\tfrac{b}{\delta}\cdot\tfrac{\beta(1-\alpha)}{\gamma+\eta+\delta}$ using $\alpha,\eta$ so as to fall into the stability regime of theorem 2.3(i).
 However, it will likely be too expensive in terms of (C3) to actually fully do this.
@@
 Finally, the bottom right displays the reproductive ratio $\mathscr{R}_{0}(s,s^{*})$ which, due to the controls, now changes with time.
-In particular, we plot the value of $\mathscr{R}_{0}(\tfrac{b}{\delta},0)$ when $\tfrac{b}{\delta}<\frac{\nu+\delta}{\overline{\mu}-\mu}$ to reflect the conditions of theorem 2.3(i) and $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$ when $\tfrac{b}{\delta}\geq\frac{\nu+\delta}{\overline{\mu}-\mu}$ to reflect the conditions of theorem 2.3(ii).
+In particular, we plot the value of $\mathscr{R}_{0}(\tfrac{b}{\delta},0)$ when $\tfrac{b}{\delta}<\frac{\nu+\delta}{\mu^{*}-\mu}$ to reflect the conditions of theorem 2.3(i) and $\mathscr{R}_{0}(\tfrac{\nu+\delta}{\mu^{*}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\mu^{*}-\mu})$ when $\tfrac{b}{\delta}\geq\frac{\nu+\delta}{\mu^{*}-\mu}$ to reflect the conditions of theorem 2.3(ii).
 In either case, if the corresponding reproductive ratio was kept below 1, the corresponding DFE would be stable, but from the plot we see that it is not optimal to do so.
 Instead, controls $\mu$ and $\nu$ are used to ensure compliance in the early stages of the epidemic, so that $\alpha$ and $\eta$ can be effectively deployed to keep the reproductive ratio low, though a small outbreak is indeed allowed to occur.
-With these cost weights, the optimal controls reduce the total cost from $12.42$ in the uncontrolled case to $2.97$ in the optimally controlled case, representing a roughly $76\%$ relative reduction in cost due to controls.
+With these cost weights, the optimal controls reduce the total cost from $8.47$ in the uncontrolled case to $4.79$ in the optimally controlled case, representing a roughly $43.5\%$ relative reduction in cost due to controls.
 Figure 2. Simulation of Scenario 1 in the absence of controls.
 Figure 3.
-Simulation of Scenario 1 with a “balanced” policy-maker: cost weights $c_{1}=1,c_{2}=0.1$.
+Simulation of Scenario 1 with a “balanced” policy-maker: cost weights $c_{1}=1,c_{2}=0.05$, $d_{1}=0.1,d_{2}=0.1$.
 The strategy is to quell noncompliance using $\mu,\nu$ so as to keep the infection rate low.
@@
 Figure 4.
-Simulation of Scenario 1 with a public health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0$.
+Simulation of Scenario 1 with a public health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0,d_{1}=0.1,d_{2}=0.1$.
 In this case, even though there is no explicit cost associated with noncompliance, because the policy-maker desires to keep infections low, the optimal strategy employs $\mu$ and $\nu$ even more strongly than in the balanced case.
@@
 We contrast this with the optimal actions corresponding to cost weights $c_{1}=5,c_{2}=0$.
-This represents a public health oriented policy-maker whose primary concern is to stop the disease, and does not even assess cost of noncompliance.
+This represents a public health oriented policy-maker whose primary concern is to stop the disease, and does not even assess cost based on noncompliance.
 The results of the simulation using these weights are in figure 4.
-In this case, the cost weight on infections is high enough that it is indeed optimal to use controls $\mu,\nu$ to ensure that, in the initial stages of the epidemic, we remain in the stability regime for the disease-free equilibrium $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$, and then use $\alpha,\eta$ to drive $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$, in essense taking advantage of theorem 2.3(i).
+In this case, the cost weight on infections is high enough that it is indeed optimal to use controls $\mu,\nu$ to ensure that, in the initial stages of the epidemic, we remain in the stability regime for the disease-free equilibrium $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$, and then use $\alpha,\eta$ to drive $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$, in essence taking advantage of theorem 2.3(i).
 This is maintained until the disease has been effectively eradicated, at which time the policy-maker eases the controls (note that, in the context of this model, the policy-maker is aware of the horizon time $T=100$, so $\alpha$ is tapered off such that infections will not significantly increase within this final time).
@@
 However, because increased noncompliance accelerates disease spread, the policy-maker must still employ $\mu,\nu$ to keep noncompliance low so as to accomplish their goal of disease elimination.
-With these cost weights, the total cost is reduced from $22.60$ in the uncontrolled case to $6.90$ in the optimally controlled case, representing a roughly $70\%$ relative reduction in cost.
+With these cost weights, the total cost is reduced from $22.60$ in the uncontrolled case to $8.62$ in the optimally controlled case, representing a roughly $61.9\%$ relative reduction in cost.
 To further explore the synergy between the controls, we include one final figure using the scenario 1 parameters, and the health oriented policy-maker with cost weights $c_{1}=5,c_{2}=0$.
@@
 While they are not displayed, the dynamics in this case look, in essence, the same as those in the top panel of figure 4 in that the disease is effectively eradicated.
-However, without access to the control $\eta$, the total cost can only be reduced from $22.60$ in the uncontrolled case to $15.43$ in the optimally controlled case, representing only a roughly $32\%$ relative cost reduction (compared to a $70\%$ relative cost reduction when $\eta$ is available).
-Thus, in some sense, controls efforts are only half as effective without $\eta$.
+However, without access to the control $\eta$, the total cost can only be reduced from $22.60$ in the uncontrolled case to $13.16$ in the optimally controlled case, representing only a roughly $41.8\%$ relative cost reduction (compared to a $61.9\%$ relative cost reduction when $\eta$ is available).
+Thus, in some sense, controls efforts are only two-thirds as effective without $\eta$.
 Figure 5.
-Simulation of Scenario 1 with a health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0$, with either treatment $\eta$ unavailable (top), or control of noncompliance $\mu,\nu$ unavailable (bottom).
+Simulation of Scenario 1 with a health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0,d_{1}=0.1,d_{2}=0.1$, with either treatment $\eta$ unavailable (top), or control of noncompliance $\mu,\nu$ unavailable (bottom).
 Note the $y$-axis change in the top left.
@@
 One sees in the bottom panel of figure 6 that without the ability to curb noncompliance, it is still beneficial to slightly decrease the reproductive ratio among the compliant population using $\alpha$ and $\eta$, but much less so since the majority of the population will become noncompliant, hampering disease prevention efforts.
-With $\mu=\nu=0$, one cannot land in the stability regime of theorem 2.3(i), nor accomplish $\mathscr{R}_{0}(\frac{\nu+\delta}{\overline{\mu}-\mu},\frac{b}{\delta}-\frac{\nu+\delta}{\overline{\mu}-\mu})<1$, so the disease is simply allowed to run its course.
-The use of $\alpha$ and $\eta$ still accomplish a reduction of total cost from $22.60$ in the uncontrolled case to $19.69$ in the optimally controlled case, representing a roughly $13\%$ relating reduction in total cost (compared to the $70\%$ cost reduction when all controls are available).
-Meaning that, with $\mu=\nu=0$, controls efforts are only about one-fifth as effective.
+With $\mu=\nu=0$, one cannot land in the stability regime of theorem 2.3(i), nor accomplish $\mathscr{R}_{0}(\frac{\nu+\delta}{\mu^{*}-\mu},\frac{b}{\delta}-\frac{\nu+\delta}{\mu^{*}-\mu})<1$, so the disease is simply allowed to run its course.
+The use of $\alpha$ and $\eta$ still accomplish a reduction of total cost from $22.60$ in the uncontrolled case to $20.31$ in the optimally controlled case, representing a roughly $10.1\%$ relative reduction in total cost (compared to the $61.9\%$ cost reduction when all controls are available).
+Meaning that, with $\mu=\nu=0$, controls efforts are only about one-sixth as effective.
+Figure 6.
+Simulation of Scenario 1 with a health oriented policy-maker: cost weights $c_{1}=5,c_{2}=0,d_{1}=0.1,d_{2}=0.1$, with either treatment $\eta$ unavailable (top), or control of noncompliance $\mu,\nu$ unavailable (bottom).
+Finally, for our last simulation in scenario 1, we demonstrate that our numerical can handle the case of $L^{1}$ control costs by setting $d_{2}=0$, eliminating the $L^{2}$ costs from (5.1).
+In this case, the Hamiltonian (3.9) is affine in $\eta,\nu\mu$, while still being quadratic in $\alpha$, since the dynamics are quadratic in $\alpha$.
+Because an affine function on an interval maximized at one of the endpoints, in this case, $\eta,\mu,\nu$ should only ever take the value 0 or their maximum value; these are akin to bang-bang controls (though they could perhaps more accurately be called “bang-no bang”).
+The control $\alpha$ does not have this structure and can still vary smoothly.
+Indeed, this is precisely what we see in figure 7.
+This demonstrates that our methods are well-equipped to handle $L^{1}$ control problems.
+It also shows why allowing for controls in $BV([0,T];U_{\text{ad}})$ is far preferable to higher regularity spaces: controls which are discontinuous but have bounded variation readily appear in applications.
+Figure 7.
+Simulation of Scenario 1 with a balanced policy-maker with no $L^{2}$ control costs: cost weights $c_{1}=1,c_{2}=0,d_{1}=0.1,d_{2}=0$.
+Theoretically, the controls $\eta,\mu,\nu$ should only take their maximum value or zero, and indeed this is borne out in the simulation.
+This demonstrates that our methods are well-equipped to handle $L^{1}$ control problems.
 5.2. Scenario 2
-In this scenario, when compared with scenario 1, we increase the disease infection rate $\beta$ and the spread rate $\overline{\mu}$ of noncompliance, and we also assume that $30\%$ of the newly introduced population members are noncompliant ($\xi=0.3$), while all other parameters are unchanged.
+In this scenario, when compared with scenario 1, we increase the disease infection rate $\beta$ and the spread rate $\mu^{*}$ of noncompliance, and we also assume that $30\%$ of the newly introduced population members are noncompliant ($\xi=0.3$), while all other parameters are unchanged.
 For this scenario, we set $c_{2}=0.01$, indicating a weak preference of the policy-maker to quell noncompliance for its own sake, though quelling noncompliance for the sake of slowing disease spread is still desirable.
@@
 When noncompliance is difficult to control, the control variables $\alpha$ and $\eta$ have a diminished effect.
-Accordingly, in this parameter regime, even a strongly public-health oriented policy-maker will not fully suppress disease spread.
-Note, for example, that the dynamics are qualitatively similar in the cases of $c_{1}=1$ and $c_{1}=6$.
-It isn’t until $c_{1}=9$ that the policy-maker significantly increases the use of the control $\alpha$ so as to more effectively stunt disease spread.
+Accordingly, in this parameter regime, only a very strongly public-health oriented policy-maker will fully suppress disease spread.
+Note, for example, that the dynamics are qualitatively similar in the cases of $c_{1}=1$ and $c_{1}=3$.
+It isn’t until $c_{1}=6$ that the policy-maker significantly increases the use of the control $\alpha$ so as to more effectively stunt disease spread.
 This indicates that in scenarios where disease spread is fast enough, and noncompliance is rampant enough, a policy-maker who is even slightly economically minded (i.e.
@@
 From top to bottom, the cost weight on infections is $c_{1}=1/3,1,3,6,9$.
-Because disease spread is so fast, suppression efforts are costly and it is not optimal to fully suppress the disease until $c_{1}$ is near 99.
+Because disease spread is so fast, suppression efforts are costly and it is not optimal to fully suppress the disease until $c_{1}\geq 6$.
 In a case like this, an economically oriented policy-maker would allow the disease to spread.
 5.3. Scenario 3
-Finally, in this scenario, we revert to the parameters of the baseline simulation in figure 3, except that we increase the spread rate for noncompliance to $\overline{\mu}=0.5$ and observe the effects of changing $\xi$, the proportion of the newly introduced population that is noncompliant.
-In particular, we have the “balanced” cost weights $c_{1}=1,c_{2}=0.1$.
+Finally, in this scenario, we revert to the parameters of the baseline simulation in figure 3, except that we increase the spread rate for noncompliance to $\mu^{*}=0.3$ and observe the effects of changing $\xi$, the proportion of the newly introduced population that is noncompliant.
+In particular, we have the “balanced” cost weights $c_{1}=1,c_{2}=0.05$.
 The full parameter list for scenario 3 is in figure 1.
@@
 A key question is then the level at which this proclivity is high enough that it is no longer optimal to attempt to suppress noncompliance.
-We include four plots in figure 7 where $\xi$ takes successive values $0,1/4,1/2,1.$ When $\xi=0$, even with the very high rate of noncompliance spread, it is optimal to fully quell noncompliance in the early stages of the epidemic.
-This is seen, for example, in the top panels of figure 7, where $\mu$ still takes its maximum value for a short time at the beginning, and noncompliance stays throughout most of the time interval (we note again that it only grows at the end because the policy-maker is aware of the horizon time, and can save some cost by releasing some controls when there is no longer time for an outbreak to occur).
-This is contrasted with the ensuing plots when $\xi=1/4,1/2,1$.
-In those cases, $\mu$ never takes it maximum value, and when $\xi=1/2,1$, $\mu$ remains below half its maximum value.
-The results for the cases $\xi=1/2$ and $\xi=1$ are very similar, seeming to indicate that for values $\xi\geq 1/2$, controls are simply ineffective.
+We include four plots in figure 9 where $\xi$ takes successive values $0,1/4,1/2,1.$ When $\xi=0$, even with the very high rate of noncompliance spread, it is optimal to significantly quell noncompliance in the early stages of the epidemic.
+This is seen, for example, in the top panels of figure 9, where $\mu$ still takes its maximum value for at the beginning, and noncompliance stays relatively low throughout most of the time interval (we note again that it only grows at the end because the policy-maker is aware of the horizon time, and can save some cost by releasing some controls when there is no longer time for an outbreak to occur).
+This behavior persists for $\xi=1/4$, but is contrasted with the ensuing plots when $\xi=1/2,1$.
+In those cases, $\mu$ is almost identically zero.
+The results int the cases $\xi=1/2$ and $\xi=1$ are very similar, seeming to indicate that for values $\xi\geq 1/2$, controls are simply ineffective.
 We note also that $\alpha$ and $\nu$ decrease with increasing $\xi$, which is a result of the fact that disease control is less effective when control of noncompliance cannot be achieved.
@@
 Simulation of Scenario 3, with increased spread rate of noncompliance, balanced cost weights, and variable proportion of the newly introduced population which is noncompliant $(\xi)$.
-Even with the increased spread of noncompliance, as long as $\xi=0$, the optimal strategy still employs the maximum available $\mu$ for a small time to suppress spread of noncompliance.
-As $\xi$ increases, slowing spread of noncompliance becomes futile, so that even when $\xi=0.25$ the policy-maker no longer uses the maximum value of $\mu$, and when $\xi=1/2$, $\mu$ is never more than half its maximum value and essentially the entire epidemic occurs on the noncompliant side.
+Even with the increased spread of noncompliance, as long as $\xi=0,1/4$, the optimal strategy still employs the maximum available $\mu$ for some time to suppress spread of noncompliance becomes futile so that when $\xi\geq 1/2$, $\mu$ is essentially zero.
 There is very little change between the simulations using $\xi=1/2$ and $\xi=1$, indicating that by the time $\xi\geq 1/2$, a threshold has been crossed and controls are ineffective.
@@
 We provide stability analysis for the disease-free equilibria of the model, and then introduce four control variables that a policy-maker may have access to: (1) reduction in infectivity corresponding to the strength of the NPIs the policy-makers enacts, (2) increase in recovery rate for compliant individuals corresponding to increased treatment efforts, (3) reduction in spread of the social contagion of noncompliance corresponding to public health information campaigns, and (4) “recovery” from noncompliance corresponding to educational campaigns aimed at increasing compliance.
-We suggest a cost functional which is quadratic in controls and jointly linear in total number of infections and total amount of noncompliance.
-We give a characterization of the optimal controls via the Pontryagin optimality principle.
+We prove existence of optimal controls for a reasonably general cost functional, which covers the case of $L^{1}$ control costs.
+We present and analyze the sequential quadratic Hamiltonian method for numerically resolving optimal controls.
 Finally, we demonstrate the behavior of our model, focusing specifically on the behavior of the optimal control maps in several parameter regimes that represent different physical scenarios and policy-maker preferences.
@@
 Acknowledgments
-CN was partially supported by the Honors Research Assistant Program (HRAP) of the University of Oklahoma.
-WW was partially supported by the Simons Foundation grant.
-Both CN and WW were partially supported by the Junior Faculty Fellowship from the University of Oklahoma.
-Appendix A Proof of Theorem 2.3
-In this appendix, we prove theorem 2.3.
-As stated in the remark below theorem 2.3, if one is willing to settle for local asymptotic stability in the case that $\mathscr{R}_{0}(s,s^{*})<1$, the theorem follows directly from [47, Theorem 2], whose hypotheses we recall and verify shortly.
-One can then upgrade to global stability using Lyapunov functions.
-We include both arguments.
-Recall, for the sake of this theorem, we are treating all control variables as constant.
-We also take $\mu\in[0,\overline{\mu})$.
-In the case the $\mu=\overline{\mu}$, the same results hold under the convention that $\overline{\mu}-\mu\sim 0^{+}$, so that $\tfrac{1}{\overline{\mu}-\mu}\sim+\infty$.
-In this case, the hypotheses of theorem 2.3(ii) cannot be met.
-Proof.
-Recall that we consider the variables $x=(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})=(I,I^{*},S,R,S^{*},R^{*})$ and the infection and transfer functions $\mathcal{F},\mathcal{V}^{+},\mathcal{V}^{-}$ defined in (2.4), (2.5).
-Defining $\mathscr{R}_{0}$ as the spectral radius of the matrix $FV^{-1}$ as described above, Theorem 2 of [47] gives local asymptotic stability of a DFE for a compartmental epidemic model when $\mathscr{R}_{0}<1$ and instability when $\mathscr{R}_{0}>1$ as long as the following hypotheses are satisfied:
-  • (H1)
-    $\mathcal{F}(x),\mathcal{V}^{+}(x),\mathcal{V}^{-}(x)\geq 0$ whenever $x$ is componentwise nonnegative.
-  • (H2)
-    For $i=1,2$, if $x_{i}=0$, then $\mathcal{V}_{i}^{-}(x)=0$.
-  • (H3)
-    For $i=3,4,5,6$, $\mathcal{F}_{i}(x)\equiv 0$.
-  • (H4)
-    If $x_{1}=x_{2}=0$, then $\mathcal{F}_{1}(x)=\mathcal{F}_{2}(x)=\mathcal{V}^{+}_{1}(x)=\mathcal{V}^{+}_{2}(x)=0.$
-  • (H5)
-    At the DFE $x$, all eigenvalues of $D\mathcal{V}^{-}(x)-D\mathcal{V}^{+}(x)$ have positive real part.
-Note that (H1)-(H4) are general hypotheses that the system must satisfy (and that our system is easily seen to satisfy), whereas (H5) is a condition specific to the DFE that one is analyzing.
-Toward verifying (H5), we see that at a DFE $x=(0,0,s,0,s^{*},0)$, $M\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=D\mathcal{V}^{-}(x)-D\mathcal{V}^{+}(x)$ has the form
-(A.1) $M=\left[\begin{smallmatrix}&&&&&\\ \gamma+\eta+\delta+(\overline{\mu}-\mu)s^{*}&-\nu&0&0&0&0\\ -(\overline{\mu}-\mu)s^{*}&\gamma+\nu+\delta&0&0&0&0\\ \beta(1-\alpha)s&(\beta(1-\alpha)+(\overline{\mu}-\mu))s&\delta+(\overline{\mu}-\mu)s^{*}&0&(\overline{\mu}-\mu)s-\nu&(\overline{\mu}-\mu)s\\ -(\gamma+\eta)&0&0&\delta+(\overline{\mu}-\mu)s^{*}&0&-\nu\\ \beta s^{*}&\beta s^{*}-(\overline{\mu}-\mu)s&-(\overline{\mu}-\mu)s^{*}&0&\nu+\delta-(\overline{\mu}-\mu)s&-(\overline{\mu}-\mu)s\\ 0&-\gamma&0&-(\overline{\mu}-\mu)s^{*}&0&\nu+\delta\\ &&&&&\end{smallmatrix}\right]$
-To find all six eigenvalues of this matrix, we note that the two eigenvalues of the first principal $2\times 2$ submatrix will also be eigenvalues of $M^{T}$ with eigenvectors of the form $v=(v_{1},v_{2},0,0,0,0)$, and thus they are also eigenvalues of $M$.
-Since $\gamma+\delta>0$, these eigenvalues are necessarily positive; this follows from Gershgorin’s theorem since the first principal $2\times 2$ submatrix is strictly (columnwise) diagonally dominant.
-The other four eigenvalues of $M$ have eigenvectors of the form $v=(0,0,v_{3},v_{4},v_{5},v_{6})$ and can be found by zooming into the last principal $4\times 4$ submatrix.
-Computing by hand, we find that these eigenvalues are
-$\lambda_{1}=\delta,\,\,\,\,\,\,\,\,\,\,\lambda_{2}=\delta,\,\,\,\,\,\,\,\,\,\,\lambda_{3}=\delta+\nu+(\overline{\mu}-\mu)s^{*},\,\,\,\,\,\,\,\,\,\,\lambda_{4}=\delta+\nu+(\overline{\mu}-\mu)(s^{*}-s).$
-These are all positive except possibly $\lambda_{4}$ so (H5) simply requires that $\lambda_{4}>0$, or
-(A.2) $s-s^{*}<\frac{\nu+\delta}{\overline{\mu}-\mu}.$
-When $\xi=0$, for $(s_{1},s^{*}_{1})=(\tfrac{b}{\delta},0)$, (A.2) reduces to $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$ which is the additional hypothesis of statement (i).
-When $\xi=0$, for $(s_{2},s_{2}^{*})=(\tfrac{\nu+\delta}{\overline{\mu}-\mu},\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu})$, (A.2) reduces to $\tfrac{b}{\delta}>\tfrac{\nu+\delta}{\overline{\mu}-\mu},$ which is the additional hypothesis of statement (ii).
-When $\xi\in(0,1]$, for $(s_{3},s_{3}^{*})$ as defined in (2.9), (A.2) reduces to
-$\tfrac{\nu+\delta}{\overline{\mu}-\mu}-\sqrt{\left(\tfrac{b}{\delta}-\tfrac{\nu+\delta}{\overline{\mu}-\mu}\right)^{2}+\tfrac{4\xi b}{\overline{\mu}-\mu}}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$
-which is always satisfied, meaning no additional hypotheses are needed.
-Thus (i), (ii), (iii) follow from [47, Theorem 2] if the statements are weakened to read “local asymptotic stability.”
-Upgrading this to global asymptotic stability is more difficult, but can be accomplished using suitably constructed Lyapunov functions as in [39], where a stochastic model similar to (2.2) is analyzed.
-So as to keep this manuscript mostly self-contained, but also refrain from repeating a very long computation, we give the construction of an explicit Lyapunov function for $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$ in the case that $\xi=0$ and $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$, and refer the reader to [39] for other cases.
-The other cases are similar, but in some cases, the construction of the Lyapunov function relies on [47, lemma 6], which says, in essence, that under assumptions (H1)-(H5), $\mathscr{R}_{0}=\rho(FV^{-1})<1$ if and only if all eigenvalues of $M\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=F-V$ have negative real part where $F,V$ are as in (2.10).
-In turn, by the theory of Sylvester equations, this implies the existence of a unique symmetric, positive definite matrix $Q$ such that $M^{t}Q+QM=-I$, and the Lyapunov function is constructed using this matrix $Q$.
-The details are in [39].
-For an explicit construction, suppose $\xi=0$ and $\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\overline{\mu}-\mu}$ so that the relevant DFE is $(s_{1},s_{1}^{*})=(\tfrac{b}{\delta},0)$.
-Let $(S,I,R,S^{*},I^{*},R^{*})$ be any solution of (2.2) and set $\Sigma=S-\tfrac{b}{\delta}$.
-Recall, solution of (2.2) are nonnegative and bounded by $\tfrac{b}{\delta}$, so $\Sigma\leq 0$ and $I,R,S^{*},I^{*},R^{*}\geq 0$.
-Writing (2.2) with this new variable $\Sigma$ (and setting $\xi=0$), we see
-(A.3) $\begin{split}\tfrac{d\Sigma}{dt}&=-\beta(1-\alpha)(\Sigma+\tfrac{b}{\delta})(I+I^{*})-(\overline{\mu}-\mu)(\Sigma+\tfrac{b}{\delta})N^{*}+\nu S^{*}-\delta u,\\ \tfrac{dI}{dt}&=\beta(1-\alpha)(\Sigma+\tfrac{b}{\delta})(I+I^{*})-(\gamma+\eta)I-(\overline{\mu}-\mu)IN^{*}+\nu I^{*}-\delta I,\\ \tfrac{dR}{dt}&=(\gamma+\eta)I-(\overline{\mu}-\mu)RN^{*}+\nu R^{*}-\delta R,\\ \tfrac{dS^{*}}{dt}&=-\beta S^{*}(I+I^{*})+(\overline{\mu}-\mu)(\Sigma+\tfrac{b}{\delta})N^{*}-\nu S^{*}-\delta S^{*},\\ \tfrac{dI^{*}}{dt}&=\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu)IN^{*}-\nu I^{*}-\delta I^{*},\\ \tfrac{dR^{*}}{dt}&=\gamma I^{*}+(\overline{\mu}-\mu)RN^{*}-\nu R^{*}-\delta R^{*}.\end{split}$
-We want to prove that the trivial solution of (A.3) is globally stable, so we define a positive definite, radially coercive Lyapunov function $V(y)$, where $y=(\Sigma,I,R,S^{*},I^{*},R^{*})$ such that $\frac{d}{dt}V(y(t))\leq-cV(y(t))$ for some positive constant $c$.
-The particular form of $V(y)$ is certainly non-unique.
-So as to cancel some nonlinear terms in the computations, we define
-(A.4) $V(y)=(\Sigma+I+S^{*}+I^{*})^{2}+A(N^{*})^{2}+B(I+I^{*})^{2}+(I^{*})^{2}+(R+R^{*})^{2}+(R^{*})^{2}$
-for some positive constants $A,B$ to be chosen later.
-Recalling that $N^{*}=S^{*}+I^{*}+R^{*}$, this is easily seen to be a positive definite quadratic form in $y$ from which we immediately have bounds $c_{1}\left|y\right|^{2}\leq V(y)\leq c_{2}\left|y\right|^{2}$ which shows that $V(y)$ is radially coercive and also reduces our goal of proving $\frac{d}{dt}V(y)\leq-cV(y)$ to proving that $\frac{d}{dt}V(y)\leq-c\left|y\right|^{2}$ for some positive constant $c.$ We compute
-(A.5) $\begin{split}\tfrac{d}{dt}V(y)&=-2\gamma(I+I^{*})(\Sigma+I+S^{*}+I^{*})-2\eta I(u+I+S^{*}+I^{*})-2\delta(u+I+S^{*}+I^{*})^{2}\\ &\hskip 28.45274pt+2A\Big{(}(\overline{\mu}-\mu)(u+\tfrac{b}{\delta}+I+R)-(\nu+\delta)\Big{)}(N^{*})^{2}\\ &\hskip 28.45274pt+2B\Big{[}\beta\Big{(}(1-\alpha)(u+\tfrac{b}{\delta})+S^{*}\Big{)}(I+I^{*})^{2}-(\delta+\gamma+\eta)I^{2}-(\delta+\gamma)(I^{*})^{2}\Big{]}\\ &\hskip 28.45274pt+2I^{*}(\beta S^{*}(I+I^{*})-\gamma I^{*}+(\overline{\mu}-\mu)IN^{*}-(\nu+\delta)I^{*})\\ &\hskip 28.45274pt+2(R+R^{*})(\gamma(I+I^{*})+\eta I-\delta(R+R^{*}))\\ &\hskip 28.45274pt+2R^{*}(\gamma I^{*}+(\overline{\mu}-\mu)RN^{*}-(\nu+\delta)R^{*})\\ &=\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}L_{1}+L_{2}+L_{3}+L_{4}+L_{5}+L_{6},\end{split}$
-where $L_{i}$ is the $i^{\text{th}}$ line of the long, preceding equation.
-We then bound line-by-line, recalling that the ultimate goal is $\frac{d}{dt}V(y)\leq-c\left|y\right|^{2}$, so we want square terms.
-We will frequently use the total population bound $\tfrac{b}{\delta}$, and the nonnegativity of each of $I,R,S^{*},I^{*},R^{*}$.
-In particular, by nonnegativity, any subset of the total population also obeys the total population bound as well.
-For example, $I+I^{*},\Sigma+\tfrac{b}{\delta}+I+R,N^{*}\leq\tfrac{b}{\delta}.$
-With this in mind, from the first line, we drop any cross terms certain to be negative (as well as some of the negative square terms that we won’t need), and then repeatedly use the Cauchy-Schwarz inequality $ab\leq\frac{\varepsilon a^{2}}{2}+\frac{b^{2}}{2\varepsilon}$ for different values of $\varepsilon>0$ to see
-(A.6) $\begin{split}L_{1}&\leq-2\delta(\Sigma^{2}+I^{2}+(S^{*})^{2}+(I^{*})^{2})-2(\gamma+\eta+2\delta)\Sigma I-2(\gamma+2\delta)\Sigma I^{*}-4\delta\Sigma S^{*}\\ &\leq-\delta\Sigma^{2}-\Big{(}2\delta-\tfrac{3(\gamma+\nu+2\delta)}{\delta}\Big{)}I^{2}+11\delta(S^{*})^{2}-\Big{(}2\delta-\tfrac{3(\gamma+2\delta)}{\delta}\Big{)}(I^{*})^{2}.\end{split}$
-For $L_{2}$, we simply use $\Sigma+\tfrac{b}{\delta}+I+R\leq\tfrac{b}{\delta},$ as well as our assumption that $\frac{b}{\delta}<\frac{\nu+\delta}{\overline{\mu}-\mu},$ to see
-(A.7) $\begin{split}L_{2}&\leq 2A\Big{(}(\overline{\mu}-\mu)\tfrac{b}{\delta}-(\nu+\delta)\Big{)}(N^{*})^{2}\leq-2\tilde{A}(N^{*})^{2}\\ &\leq-2\tilde{A}((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2}).\end{split}$
-where $\tilde{A}=A((\nu+\delta)-\tfrac{b}{\delta}(\overline{\mu}-\mu))>0$ can be made arbitrarily large by taking $A$ large.
-Continuing with $L_{3}$, the difficulty is how to handle the triple product term $\Big{(}(1-\alpha)(\Sigma+\tfrac{b}{\delta})+S^{*}\Big{)}(I+I^{*})^{2}$, bearing in mind that we want to keep the full effect of the $(1-\alpha)$ because this appears in $\mathscr{R}_{0}(\tfrac{b}{\delta},0).$ Thus we split up the $(\Sigma+\tfrac{b}{\delta})$ and the $S^{*}$ in the first term and use
-(A.8) $\begin{split}(\Sigma+\tfrac{b}{\delta})(I+I^{*})^{2}&\leq(\tfrac{b}{\delta})I^{2}+2(\tfrac{b}{\delta})II^{*}+(\tfrac{b}{\delta})(I^{*})^{2}\\ &\leq\Big{(}(\tfrac{b}{\delta})+\varepsilon\Big{)}I^{2}+C\Big{(}1+\tfrac{1}{\varepsilon}\Big{)}(I^{*})^{2}\end{split}$
-where $\varepsilon>0$ will be specified later, and $C$ is some positive constant comprised of the other parameters (in what follows, this constant may change from line-to-line but it will always be a finite, positive constant).
-Next, we use
-(A.9) $\begin{split}S^{*}(I+I^{*})^{2}&\leq(\tfrac{b}{\delta})S^{*}(I+I^{*})\\ &\leq\varepsilon I^{2}+C(1+\tfrac{1}{\varepsilon})(S^{*})^{2}+C(I^{*})^{2}.\end{split}$
-Combining (A.8) and (A.9) with the other terms in line 3 of (A.5), we have
-(A.10) $L_{3}\leq 2B\Big{[}(1-\alpha)\Big{(}(\tfrac{b}{\delta})+2\varepsilon)-(\gamma+\delta+\eta)\Big{]}I^{2}+C(1+\tfrac{1}{\varepsilon})((S^{*})^{2}+(I^{*})^{2}).$
-For line 4 in (A.5), we use $I,I+I^{*}\leq\tfrac{b}{\delta}$ so that all terms are simply quadratic in $S^{*},I^{*},R^{*}$ whereupon
-(A.11) $L_{4}\leq C((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2}).$
-For line 5 in (A.5), the only bothersome terms are those involving $I$ and $R$, and we have
-(A.12) $L_{5}\leq\tfrac{2(\gamma+\eta)}{\delta}I^{2}-\delta R^{2}+C((I^{*})^{2}+(R^{*})^{2}).$
-And finally, for line 6 of (A.5), once we bound $R\leq\tfrac{b}{\delta}$, all terms are quadratic in $S^{*},I^{*},R^{*}$ again so
-(A.13) $L_{6}\leq C((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2}).$
-Finally, combining the six bounds (A.6), (A.7),(A.10),(A.11),(A.12),(A.13), we arrive at
-(A.14) $\begin{split}\tfrac{d}{dt}V(y)&\leq-\Big{(}2\delta-2B\Big{[}(1-\alpha)\Big{(}(\tfrac{b}{\delta})+2\varepsilon)-(\gamma+\delta+\eta)\Big{]}-\tfrac{3(\gamma+\nu+2\delta)+2(\gamma+\eta)}{\delta}\Big{)}I^{2}\\ &\hskip 28.45274pt-\delta(\Sigma^{2}+R^{2})+\Big{[}C\Big{(}1+\tfrac{1}{\varepsilon}\Big{)}-2\tilde{A}\Big{]}((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2})\\ &=\Big{(}2\delta+\tfrac{2B}{\gamma+\eta+\delta}\Big{[}1-\mathscr{R}_{0}(\tfrac{b}{\delta},0)-\tfrac{2\varepsilon}{\gamma+\eta+\delta}\Big{]}-\tfrac{3(\gamma+\nu+2\delta)+2(\gamma+\eta)}{\delta}\Big{)}I^{2}\\ &\hskip 28.45274pt-\delta(\Sigma^{2}+R^{2})+\Big{[}C\Big{(}1+\tfrac{1}{\varepsilon}\Big{)}-2\tilde{A}\Big{]}((S^{*})^{2}+(I^{*})^{2}+(R^{*})^{2}).\end{split}$
-Since we have assumed $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<1$, we can take $\varepsilon>0$ small enough that $1-\mathscr{R}_{0}(\tfrac{b}{\delta},0)-\tfrac{2\varepsilon}{\gamma+\eta+\delta}>0$, and then we can choose $B>0$ so that
-$\tfrac{2B}{\gamma+\eta+\delta}\Big{[}1-\mathscr{R}_{0}(\tfrac{b}{\delta},0)-\tfrac{2\varepsilon}{\gamma+\eta+\delta}\Big{]}-\tfrac{3(\gamma+\nu+2\delta)+2(\gamma+\eta)}{\delta}=0.$
-Finally, having fixed $\varepsilon>0$, we simply take $A$ large enough so that $C(1+\tfrac{1}{\varepsilon})-2\tilde{A}<0$, and we arrive at $\tfrac{d}{dt}V(y)\leq-c\left|y\right|^{2}$ for a positive constant $c$, whereupon the DFE $(s,s^{*})=(\tfrac{b}{\delta},0)$ is globally asymptotically stable.
-As stated before, we refer the reader to [39] for other cases. ∎
-Chloe Ngo, Honors College, University of Oklahoma, Norman, OK, USA; e-mail: chloe.ngo@ou.edu
-Christian Parkinson, Department of Mathematics & Department of Computational Mathematics, Science and Engineering, Michigan State University, East Lansing, MI, USA; e-mail: chparkin@msu.edu
-Weinan Wang, Department of Mathematics, University of Oklahoma, Norman, OK, USA; e-mail: ww@ou.edu
+Chloe Ngo was partially supported by the Honors Research Assistant Program (HRAP) of the University of Oklahoma.
+Weinan Wang was partially supported by the Simons Foundation grant.
+Both Chloe Ngo and Weinan Wang were partially supported by the Junior Faculty Fellowship from the University of Oklahoma.